T.R | Title | User | Personal Name | Date | Lines |
---|
719.1 | Weyers Appointed Goddard Director of Flight Projects | PRAGMA::GRIFFIN | Dave Griffin | Mon Jul 08 1991 21:29 | 18 |
| Release No. 91-121 (7/3/91)
Vernon J. Weyers, Director of Space Flight Systems at the Lewis
Research Center, Cleveland, OH, will be appointed as Director of Flight
Projects, Goddard Space Flight Center, Greenbelt, MD, effective September 1,
1991, Goddard Director Dr. John M. Klineberg announced today. Weyers replaces
Peter T. Burr who was reassigned to the position of deputy director of the
Goddard Space Flight Center in February 1991. Weyers began his career at the
Lewis Research Center in 1962 supporting the Apollo Program and subsequently
the Viking and Voyager missions. He served as project manager of a joint
turbine development effort with the Department of Interior; Lewis Institutional
Manager of the Shuttle/Centaur Project Office; and Chief of the Systems
Engineering and Integration Division in the Space Station Freedom Directorate.
Weyers has served as Director of Space Flight Systems at the Lewis Research
Center since January 1988. Weyers earned his B.S. degree in Electrical
Engineering from the University of Wisconsin in 1962 and his M.S. degree in
Electrical Engineering at the University of Toledo in 1970. He received the
NASA Exceptional Service Medal in August 1989.
|
719.2 | GSFC Project Status for April 1993 | PRAGMA::GRIFFIN | Dave Griffin | Mon Apr 12 1993 14:31 | 89 |
| Highlights:
Astro-D
Compton Gamma Ray Observatory (Compton)
Cosmic Background Explorer (COBE)
Extreme Ultraviolet Explorer (EUVE)
Hubble Space Telescope (HST)
Solar Anomalous and Magnetospheric Particle Explorer (SAMPEX)
Upper Atmosphere Research Satellite (UARS)
Science Note: SN 1993J
Astro-D -- The instruments on the Astro-D spacecraft launched in Japan on
February 20, 1993 are being activated. The two Japanese focal plane
instruments, the Gas Scintillation Imaging Spectrometers (GISs), were turned on
in mid- March. Both GIS counters have been verified as working normally.
GSFC's Solid State Imaging Spectrometer (SIS) Charge Coupled Device (CCD)
cameras also are being activated. Both camera vent valves have opened and the
cameras currently are stabilized around -81 degrees Farenheit (-63 degrees
Centigrade).
Compton Gamma-Ray Observatory -- The first delivery of public Energetic
Gamma-Ray Experiment Telescope (EGRET) data into the archive has been made.
This data comprises photon lists, exposure histories and maps from the
verification period through July 26, 1991. As of March 25, the mean orbital
altitude was 233.11 statute miles (359.06 kilometers). Planning and rehearsal
simulations for the orbit reboost operations continue. The series of daily
orbit reboost burns are planned to begin June 15 and continue for about two
weeks.
Cosmic Background Explorer -- COBE continues to acquire all science and
engineering data without any major problems or operational errors. The COBE
Operations team is still attempting to firm up contingency plans in the event
of a gyro failure during the upcoming eclipse season, which occurs from May
through July.
Extreme Ultraviolet Explorer -- Observations of the planet Jupiter and the
stars DELTA-CEN and HD-131156 were planned for the first week in April,
followed by a short observation of the moon. The next sky map "gap filling"
period is scheduled for April 19 through 25.
Hubble Space Telescope -- On Wednesday, March 24, the HST spacecraft entered
Software Sun Point Safe Mode. The preliminary indication was that the solar
array electronics had sensed an erroneous position for a solar array and
transmitted this information to the onboard computer, which in turn responded
by safing the spacecraft systems. This initial level of safing suspends
science operations but retains all vehicle control through the operations
flight computer (DF-224). Initial analysis of the telemetry indicated that the
safing had been caused by a miscompare between the actual position of the solar
arrays and the position indicated by the Solar Array Drive Electronics (SADE)
resolver. The most likely cause is believed to be either a failure in the DC
power supply or the resolver electronics. The SADE are fully redundant and
operations have been switched to the correctly operating side two. Evaluation
and analysis is continuing and science operations have been resumed. The
number of observations may be reduced by about 20 percent for the next several
weeks until flight software safing modes can be modified. Once this is
accomplished, normal science operations can resume. In the event of a
subsequent failure of side two (SADE 2), science observations could still
continue by factoring power requirements into the constraints during the
planning and scheduling of the observations.
Solar Anomalous and Magnetospheric Particle Explorer -- SAMPEX acquired all
science and engineering data without any significant problems or operational
errors. The spacecraft continues to perform well while it is in continuous
sunlight.
Upper Atmosphere Research Satellite -- UARS, launched September 12, 1992,
surpassed its 18-month mission success criteria mark. More than 100 scientists
participated in a UARS workshop in Virginia Beach, Va., last month,
commemorating this occasion. More than 55 studies were presented. The
scientific focus of the workshop was high latitude processes affecting ozone in
the stratosphere and mesophere although one session highlighted higher altitude
phenomena and wind measurements. Many of the talks illustrated new
understanding of the development and distributions of important gaseous species
impacting ozone, and a newly perceived role of dynamics and transport from high
altitudes in the formation and maintenance of the Antarctic "ozone hole."
Science Note: A supernova discovered on March 30, 1993, is the target of
opportunity for three satellites. Observed initially by NASA's International
Ultraviolet Explorer, SN 1993J also has been looked at in the gamma-ray
spectrum by NASA's Compton Gamma-Ray Observatory and in the x-ray spectrum by
the Japanese Astro-D satellite and the German Roentgen Satellite.
|
719.3 | Satellite Updates for January 6-12 | VERGA::KLAES | Quo vadimus? | Thu Jan 13 1994 17:37 | 340 |
| Article: 5212
Newsgroups: sci.space.news
From: Steve Andrews <[email protected]>
Subject: Orbiting Satellites
Sender: [email protected]
Organization: NASA/GSFC Guidance and Control Branch
Date: Wed, 12 Jan 1994 16:30:04 GMT
Orbiting Satellites
For the week of January 6 to January 12, 1994
From Goddard Space Flight Center
Code 602 Project Manager Paul J. Pashby
EUVE, ICE, IMP, IUE,
ERBS, COBE, SAMPEX, GRO
Since this news release was published at the beginning of the
holiday season, I'm sure many of you might have not seen it.
Because it explains in good detail what was really involved in the
GRO reboost manuever, I thought it was appropriate to publish it
again for those who may have missed it the first time around.
Michael Finneran Dec. 20, 1993
Goddard Space Flight Center, Greenbelt, Md.
(Phone: 301/286-5565)
RELEASE NO. 93-159
NASA SUCCEEDS WITH GAMMA-RAY OBSERVATORY REBOOST
NASA has successfully boosted its Compton Gamma-Ray Observatory
into a higher orbit around the Earth, a move that will keep the
17-ton (15.4 metric ton) satellite from reentering the atmosphere,
said officials at Goddard Space Flight Center in Greenbelt, Md.,
which manages the spacecraft.
The reboost is a significant accomplishment because it extends the
mission life of the observatory by 5 years and prevents a reentry
in which large parts from the spacecraft could have struck Earth.
The reboost, which concluded Dec. 17, had been anticipated when the
2-year-old spacecraft was designed. It was necessary because of
solar activity that slowly pushed the satellite back toward Earth.
Initially scheduled earlier in the mission, the reboost was delayed
because of the high risk of operating the spacecraft's propulsion
system after a problem that occurred shortly after launch.
"Although this reboost always was planned, its successful
completion involved solving some complex problems never before
encountered by NASA," said Goddard's Tom LaVigna, former deputy
project manager for the observatory, who retained management
responsibility for the reboost.
NASA managers had to compensate for a balky rocket thruster on the
observatory, LaVigna said, and devise a new way to inject fuel into
the sensitive plumbing lines leading to the thrusters.
The complex procedure for boosting Compton was developed by Goddard
engineers after extensive analysis and simulations. It employed a
novel way to operate the observatory propulsion and the attitude
control subsystems together to achieve a controlled boost.
The observatory has the largest propulsion system ever launched by
NASA, carrying 4,200 pounds (1,910 kilograms) of propellant. It has
eight attitude control thrusters and four larger orbit adjust
thrusters, with half of each split between the "A" and "B" sides of
the spacecraft.
Through mathematical analyses and tests, it was determined that
very high fuel-line pressure surges had compromised the A-side
propulsion system when the system first was operated shortly after
launch. After extensive study and test, a unique fuel-line priming
procedure was developed to prepare the redundant B-side thrusters
for use in the reboost.
The priming worked well, but the reboost was terminated when
the B2 attitude control thruster performed significantly below
the other thrusters and caused the observatory to go unstable and
tumble. Quick action by ground controllers terminated the thruster
firings and, using a contingency procedure, brought the observatory
back into a stable mode.
With the standard procedure no longer usable, Goddard engineers
developed a novel way around the problem. The procedure involved
firing two of the four 100-pound (45.4-kilogram) orbit adjust
thrusters for 60 to 90 seconds. The 5-pound (2.3-kilogram) attitude
control thrusters and reaction wheels were used together to limit
unstable conditions and errors in pointing the observatory. A total
of 18 propulsion system burns were used, one each day.
The reboost was accomplished in two phases, one that concluded in
October and another spanning November and December. The first
phase lifted the observatory's apogee -- or highest point in orbit
-- to 280 miles (452 kilometers). The apogee had descended to 214
miles (346 kilometers). The two-part second phase of the reboost
nudged the spacecraft's perigee -- lowest point -- to the same 280
miles (452 kilometers), making Compton's orbit around the Earth
almost perfectly circular.
If the satellite had descended below 180 miles (290 kilometers), it
could have lost the stability required for reboost, LaVigna said.
If the observatory was not reboosted, he added, it would have
dropped to that altitude in late April of 1994.
Two instruments on the spacecraft would not have burned up entirely
if they had reentered the atmosphere -- the Energetic Gamma-Ray
Experiment Telescope (EGRET) and the Compton Telescope (COMPTEL).
Each have very heavy, dense, sodium-iodide gamma-ray detectors.
The EGRET detector is 2.3 feet (0.7 meters) square by eight inches
(20 centimeters) thick and weighs 900 pounds (408 kilograms). The
COMPTEL detector is 3.4 feet (1.04 meters) in diameter and two
inches (five centimeters) thick and weighs 300 pounds (136
kilograms).
When the Gamma-Ray Observatory's life is over, NASA managers plan
a controlled reentry so that the unburned debris falls harmlessly
on an uninhabited area.
The $550-million observatory, a collaboration between the United
States, Germany and The Netherlands, has a mission goal of two to
six years and carries four science-gathering instruments. The
satellite exceeded its minimum life in April this year. The
observatory was sent aloft to study the gamma-ray universe. Gamma
rays are the highest energy radiations in the electromagnetic
spectrum, created by some of the most violent events in the
universe, such as the explosion of stars. Gamma rays can be studied
effectively only from space because Earth's atmosphere prevents
them from reaching the ground.
The Compton Gamma-Ray Observatory is managed by Goddard for the
Office of Space Science at NASA Headquarters in Washington, D.C.
The observatory was launched April 5, 1991, on the Space Shuttle
Atlantis and is the heaviest civilian spacecraft ever deployed by
a shuttle.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GAL 304-01 science target on January 3.
The next maneuver is planned for January 16 as an adjustment to the
current attitude.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of
full-time coverage achieved with 32 Kbps real-time data from
January 3 through January 9 was 79 percent, including 9 percent
from the GRO Remote Terminal System. An additional 8 percent
coverage was achieved with the BATSE 1 Kbps subset of data during
this time.
Operational testing for engineering evaluation of the GRO Remote
Terminal System (GRTS) is continuing very successfully. As TDRS-1
(TDZ) is drifting closer to its final destination, the GRTS
coverage is increasing. TDRS-1 is expected to reach 85 degrees
East longitude in early February 1994. When GRTS becomes
operational in February, it will provide a significant increase in
real-time coverage for GRO.
The performance of the batteries continues to be monitored very
closely, with power configurations adjusted as necessary. The
observatory is functioning normally with primary systems powered
from modular power subsystem 2, which is performing very well.
Modular power subsystem 1 continues to support a reduced load of
mainly heaters.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis. During
the week of 01/03/94 through 01/09/94, 2 real time passes were
supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 01/01/94 through 01/09/94, 7 real time passes
were supported with 29 commands transmitted.
ICE/IMP TPOCC - G. Bullock
String 1 was upgraded from a 380 HP system to a 715 HP system on
January 5, 1994. The test team will begin testing January 10,
1994.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
Several observations of comet Enke were performed this week. The
sun angle for these observations ranged from Beta = 122 Degrees to
Beta = 126 Degrees. The observations performed at the higher end
of this range were power negative. During these observations the
Santiago and Vilspa stations provided backup support, however all
operations were nominal during these observations.
On 1/5/94 at 2122 Z the Sigma 5 computer, being used to support
real time operations, crashed and could not be rebooted. The
backup computer was brought up and used for support during this
time. After being reloaded, the Sigma 5 computer was brought back
on line at 0015 Z on 1/6/94. No loss of science time resulted due
to this event because it occurred during a long exposure.
On 1/8/94 the normal command and telemetry link between the
Wallops station and the IUEOCC was down for approximately six
hours while a scheduled reconfiguration of the lines occurred; the
DOMSAT satellite was replaced. During this period the TDMA
telemetry link and the DDPS command system were used. No
problems were experienced with the backup modes of support and
operations were not interrupted.
Several instances of the intermittent problems recently being
experienced with the primary 8391 telemetry line and the primary
8430 command line were experienced this week. During these
occurrences the telemetry data stream was rerouted through the
TDMA line and command support was provided through the DDPS
system. Only five minutes of science time was lost as a result of
these line problems. For the most part, these problems occurred at
times after a science observation had been completely setup and
initiated, therefore the science experiment being performed by the
spacecraft continued despite the problems experienced with the
ground system.
OPERATIONS
A total of 91 images were taken during the week:
Long Wavelength Primary (LWP). . . . . . . . 46
Long Wavelength Redundant (LWR). . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . 45
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .5 minute
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6715 minutes
Effectiveness. . . . . . . . . . . . . . . .99.9%
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
EUVE continues to perform very well. No new spacecraft or payload
problems to report. The Target of Opportunity U-Geminorum was
viewed several times last week in an effort to get data on the
declining side of the star's light curve. Early results indicate
that this was accomplished.
EUVE is currently observing the quasar 3C273 until day 14 when it
will slew to Geminga for a ten day observation.
SPACECRAFT STATUS
EUVE's Explorer Platform continued its overall good performance.
An intermittent guide star acquisition problem with FHST #2
occurred during the observation of WD-1057 (DOY 7). Preliminary
analysis indicates the anomalous behavior was caused by other
(blocker?) stars within the field of view or near the edge of the
tracker. Additional data analysis is being performed by GSFC's
Flight Dynamics Facility. All health and safety telemetry from the
spacecraft was in the nominal range. An operator error on day 7
placed both tape recorders briefly in the fast forward mode, but
real time data was being received at the time by UCB and so there
was no science data loss.
A slightly modified Telemetry Monitor (TMON) group designed to
improve the trigger point of the spacecraft into constant current
mode based upon battery state of charge was successfully uplinked
on day 3. It is currently being evaluated for its stabilizing
effect on battery performance.
INSTRUMENT STATUS
Scanner C and detector #7 were turned off for the observation of
U-Gem. Scanner C was turned off so that more data could be
processed from the spectrometers, while detector #7 was turned off
because the EUV count rate from U-Gem was so high that there was
concern at UCB about damage to the detector.
The focal plane thermistors for detectors #2 and #3 also went in
and out of high yellow limit during this observation, but returned
to normal after slewing to the next target.
UPCOMING EVENTS
EUVE will observe 3C273 until 014/2121Z when it begins a ten day
observation of Geminga.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecaft is still operating with only one battery with
two failed cells, while acquiring scientific data from both the
ERBE-NS and the SAGE instruments. The ERBE-NS instrument will be
turned OFF today in preparation for the spacecraft yaw-turn which
will be performed later this week. This maneuver must be performed
every 4-5 weeks in order to keep sunlight on the solar arrays. It
is performed during maximum eclipses, and without any feedback
control since most of the spacecraft gyros have failed. The timing
on these thruster commands is very critical and the resulting ACS
errors are also very sensitive to the battery voltage. The Lead
ERBS Control Center Analyst broke his leg on the ice outside
Building #14 this week.
COSMIC BACKGROUND EXPLORER (COBE) - J. Williamson
The science data acquisition on COBE has been terminated and the
Control Center is now conducting End-of-Life tests on the
spacecraft and the instruments. After all of these tests are
completed in approximately two weeks, the spacecraft will be
transferred to Wallops (Code 822) and to the STGT Project (Code
530) for them to use as a test satellite for their respective
projects. The needed spacecraft operating instructions and a
highly condensed list of spacecraft commands have already been
supplied to these projects, and their spacecraft testing will
start in about two weeks. After this phase has been completed all
COBE support will be terminated at GSFC.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
spacecraft operations have moved back to their regular Mission
Operations Room (MOR); however, additional hardware and software
upgrades are needed before the MOR reaches the configuration
needed for the FAST launch. In the next few weeks, SAMPEX will be
testing a number of FAST upgrades at the various ground stations.
Steve A.
[email protected]
|
719.4 | January 13-27 | VERGA::KLAES | Quo vadimus? | Thu Jan 27 1994 11:53 | 274 |
| From: US1RMC::"[email protected]" "Steve Andrews" 27-JAN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 1/13/94
Orbiting Satellites Jan 13 to Jan 19, 1994
From Goddard Weekly Report
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis has
been oriented toward the GAL 304-01 science target since January 3.
A small adjustment to the attitude was made on January 16. The next
attitude maneuver is planned for January 23 to the CEN A target.
Operational testing for engineering evaluation of the GRO Remote
Terminal System (GRTS) is continuing very successfully. As TDRS-1
(TDZ) is drifting closer to its final destination, the GRTS coverage
is increasing. The percentage of full-time coverage achieved with 32
Kbps via GRTS from January 10 through January 16 was 13 percent. TDRS-1
is expected to reach 85 degrees East longitude in early February 1994.
The performance of the batteries continues to be monitored very
closely, with power configurations adjusted as necessary. The
observatory is functioning normally with primary systems powered from
modular power subsystem 2, which is performing very well. Modular
power subsystem 1 contineus to support a reduced load of mainly heaters.
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft is still operating with only one battery with two
failed cells, while acquiring scientific data from the SAGE
instrument. Another spacecraft 180 degree yaw turn was successfully
completed last week. This maneuver, which is required to keep
sunlight on the solar arrays, is performed every 4-5 weeks during
deepest eclipses. The ERBE-NS instrument will be powered ON this week
after being OFF for eight days during the maximum eclipses. The
Project again thanks the Code 734 Power Engineers and the ERBS Control
Center (Code 519.1/ATSC) personnel for their excellent job on keeping
the ERBS mission going and acquiring useful data despite the highly
degraded battery.
COSMIC BACKGROUND EXPLORER (COBE)
All science data acquisition on COBE was terminated on 12/23/93, and
almost all of the End-of-Life Tests were completed at the end of last
week. The Flight Ops Team are continuing minor tests this week, and
they will verify the ability of Wallops to command the COBE spacecraft
transponder ON/OFF this week.
The COBE spacecraft will then be turned over to Wallops and to the
STGT Project next week for them to use COBE as a communications test
satellite for their respective projects. The STGT interface tests
with the COBE spacecraft will be performed in a few weeks after STGT
completes the installation of their ground equipment at their remote site.
[Author's Note: STGT = Second TDRS Ground Terminal, if my memory is correct]
SAMPEX
SAMPEX continues to acquire all science and engineering data without
any significant problems or operational errors. All of the FAST
Operational testing was moved into the SMEX [Small Explorer] Control
Center last week, after Code 513 [Project Operations Branch]
determined that there was not enough space in the SOTA (Spacecraft
Operations and Tests Area) for both FAST Ops testing and the software
testing of newer missions, e.g., XTE. This presents an operational
problem since the SMEX Control Center is still being upgraded with new
hardware, software, and other equipment.
Steve A.
[email protected]
From: US1RMC::"[email protected]" "Steve Andrews" 27-JAN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 1/20/94
Jan 20 Jan 26 94
Orbiting Satellites for 1/20/94 to 1/27/94
EUVE, ICE, IMP, IUE,
ERBS, COBE, SAMPEX, GRO
International Cometary Explorer (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 01/10/94 through 01/16/94, 2 real time passes were
supported with no commands transmitted.
Interplanetary Monitoring Platform (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 01/10/94 through 01/16/94, 7 real time passes
were supported with 30 commands transmitted.
A Geospace Environment Modeling (GEM) campaign started on January
11 at 1600Z and ended on January 14 at 1600Z. Results of data
recovered from IMP-8 during this period will be computed when
reports are received from all stations.
International Ultraviolet Explorer (IUE)
HIGHLIGHTS
On January 10, patches to the onboard computer software needed for
the control modes used during an eclipse were put in line in
preparation for shadow season #33. These patches are put in line
just prior to the onset of a shadow season and then taken out
following the completion of the shadow season. Shadow season #33
will run from January 11 through February 8. The insertion of these
patches into the onboard computer code was performed without incident.
Shadow season #33 began on January 11. To date the maximum depths
of discharge for battery #1 and battery #2 have been 31.25 % and
42.61 % respectively. So far, the spacecraft has performed as
expected during this shadow season. The maximum daily eclipse will
occur on January 25 with an umbra of approximately 59 minutes. On
January 21, the cycling off and on of VHF telemetry during the
umbra portion of each daily eclipse was initiated. The telemetry
system is cycled off and on with a five minute period, four minutes
with no telemetry and one minute with telemetry. This action is
taken in order to reduce the depth of discharge on the batteries.
VILSPA supported operations may be affected by a general strike
throughout Spain on January 27. In preparation for this event,
backup support required for the daily eclipse period has been
requested through the Santiago station. In the event that VILSPA
is unable to assume control of the spacecraft during their
regularly scheduled period on this day, operations will continue
from the GSFC control and science centers.
On January 21 and 22, two observations at the low Beta angles of 28
and 29 degrees were successfully performed. VHF telemetry support
was scheduled for these observations in order to reduce the power
load; however, VHF telemetry was not needed since the spacecraft
was power positive at these sun angles.
Several power hits occurred throughout the week requiring the ground
system to be rebooted; however, there was minimal impact to operations.
A Delta-V is planned for the week of February 14. FDF is still
working on the final details for this orbit adjustment maneuver.
OPERATIONS
A total of 81 images were taken during the week of 01/10.94
through 01/16/94.
Long Wavelength Primary (LWP) . . . . . . . . 42
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . . 39
A total of 74 images were taken during the week of 01/17.94
through 01/23/94.
Long Wavelength Primary (LWP) . . . . . . . . 34
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . . 40
OPERATIONAL EFFECTIVENESS
Operational effectiveness for the week of 01/10/94 through
01/16/94 was:
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minute
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
Operational effectiveness for the week of 01/17/94 through
01/23/94 was:
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering Tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . 100 %
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
NASA's Extreme Ultraviolet Explorer, launched on June 7, 1992,
recently passed the orbit-9000 mark. The spacecraft bus and EUVE
science payload continue to perform remarkably well. This week's
telescope time was devoted to the observation of Geminga. An
additional weather related outage at GSFC PACOR caused a temporary
loss of realtime data to UCB. All service has now been restored.
SPACECRAFT STATUS
The Explorer Platform continued its overall good performance this
week. No out of limit telemetry values for any subsystem were
noted. Battery performance has now stabilized very nicely with the
careful management of inrush current, temperature, and total
charge. Autonomous solar array movements to maintain power and
avoid dangerous solar reflection from the s/c body onto array #1
have been proceeding smoothly. Occasionally, due to the effect of
orbit precession during long observations, ground commands are sent
to the arrays to fine tune their inclination to the sun. Such a
command was necessary this week. The High Gain Antenna (HGA)
gimbal and Steering Control Electronics (SCE) temperatures managed
by the manipulation of the HGA shadow- continue to be monitored
closely to guard against any reoccurrence of the break-track
incident during July '92. Gyro channel B is still performing at
elevated noise levels but with no adverse effect on the platform's
attitude stability. All other subsystems (C&DH, Power/thermal,
attitude control) are performing very well, occasionally
manifesting only (so far!) minor anomalies.
A statewide power emergency in Maryland on the 20th resulted in the
closure of GSFC to all non-essential personnel and the temporary
outage of the PACOR facility. This caused a loss of realtime
telemetry to UCB. At the Project's request, power was restored for
a one-string, realtime-only operation in PACOR. Production data
was backlogged until the 14th.
INSTRUMENT STATUS
No new problems or incidents to report.
UPCOMING EVENTS
EUVE will observe Geminga until the 24th followed by observations
of RE 1309, Spica, RE 2127-221 and RE 1016-053.
EARTH RADIATION BUDGET SATELLITES (ERBS) - J. Williamson
The ERBS spacecraft is still operating with only one battery with
two failed cells, while acquiring scientific data from both the
SAGE and ERBE-NS instruments. The spacecraft will enter a 10-day
period of continuous sunlight in about one week.
COSMIC BACKGROUND EXPLORER (COBE) - J. Williamson
All science data acquisition on COBE was terminated on 12/23/93,
and all of the End-of-Life Tests were completed at the end of last
week. The Flight Ops Team is now completing the off-line analysis
for these tests, and clearing out the Control Center. The COBE
spacecraft is now the responsibility of Wallops and the STGT
Projects, who will use COBE as a communications test satellite for
their respective projects.
The Wallops station has completed their interface tests with the COBE
spacecraft several weeks ago, and similar tests will be performed with
the STGT station in about two weeks. The COBE Project has received an
additional request to use COBE over Antarctica from Code 450 [TOMS Project
- sfa]. The Project has referred this request to Wallops (Code 822).
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. New SAMPEX
software is being tested in the SOTA (Spacecraft Operations and
Test Area), and the SMEX Control Center is still awaiting new
hardware, and other items. In addition to the regular SAMPEX flight
Ops, a FAST Mission Readiness Test will also be run this week from
the SMEX Control Center.
Steve A.
[email protected]
|
719.5 | January 27 to February 2 | VERGA::KLAES | Quo vadimus? | Mon Feb 07 1994 18:59 | 186 |
| From: US1RMC::"[email protected]" "Steve Andrews" 5-FEB-1994
To: [email protected]
CC:
Subj: Orbiting Satellites - 1/27/94
From the Goddard Weekly Report
Code 602 Orbiting Satellites Paul J. Pashby
For the week: Jan. 27, 1994 to Feb. 2 1994
EUVE, ICE, IMP, IUE,
ERBS, COBE, SAMPEX, GRO
EXTREME ULTRAVIOLET EXPLORER (EUVE)
The EUVE satellite conducted a routine week of operations. Four
stellar targets and the Moon were observed. All payload and
spacecraft bus subsystems are operating nominally. Three targets
are scheduled for the coming week.
SPACECRAFT STATUS
EUVE's Explorer Platform continued to perform well this week. Two
TDRSS passes experienced some difficulty on day 28 - this due to a
miscue between the Flight Operations Team and the GSFC Flight
Dynamics Facility over the most current FDF planning products.
All the realtime data lost during these events was later recovered
from tape recorder playbacks. An unexpected OBC status buffer
message (SBM 704) on day 23 was traced to a minor flight software
bug associated with an incrementing counter. No corrective action
is planned at this time.
The backlog of production science data due to the recent PACOR
weather-related outages is now gone. Realtime and production data
are being produced on time.
INSTRUMENT STATUS
Detectors 1,2,3 and 7 were turned off for the observation of the
Moon. The rate shutdown parameters were also adjusted so that more
lunar photons would be processed. No other problems or incidents
to report.
UPCOMING EVENTS
EUVE will observe RE 1016-053, PSR J0437 and VELA in the coming week.
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft is still operating with only one battery with
two failed cells, while acquiring scientific data from the ERBE-NS
instrument. The SAGE science data has been suspended for 10-days
while the spacecraft is in continuous sunlight.
Special battery charge control procedures are used while the
spacecraft is in continuous sunlight in order to keep from over
charging the battery.
COSMIC BACKGROUND EXPLORER (COBE)
The COBE Mission was terminated last week. All of the End-of-Life
tests have been completed, and the COBE Control Center has gone out
of business. Some of the Flight Ops Team will remain another two
months writing End-of-Life reports and closing down of the Project.
The COBE spacecraft has been transferred to TOTS Project at Wallops
and to STGT Project here at GSFC for them to use as a communications
test satellite for their ground stations.
SAMPEX
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. New SAMPEX
hardware upgrades are expected to be delivered to the Control
Center this week, and new SAMPEX software is expected to be
delivered early next week. SAMPEX shares the SMEX Control Center
with FAST which is scheduled to be launched in about 6 months. The
Control Center is already quite busy with all the changes, testing
and the SAMPEX operations that are all occurring simultaneously.
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the CYG X1 instead of the previously scheduled
GAL 015+06 science target on February 1. The CYG X1 was declared
a target of opportunity and will be observed for one week. The next
attitude maneuver is planned for February 8 to the GAL 357-11 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of
full-time coverage achieved with 32 Kbps real-time data from
January 23 through January 28 was 79 percent, including 17 percent
from the GRO Remote Terminal System. An additional 7 percent
coverage was achieved with the BATSE 1 Kbps subset of data during
this time.
Operational testing for engineering evaluation of the GRO Remote
Terminal System (GRTS) is continuing very successfully. As TDRS-1
(TDZ) is drifting closer to its final destination, the GRTS
coverage is increasing. TDRS-1 is expected to be on station at 85
degrees East longitude on February 7, 1994. Current plans are to
suspend acquisition of GRO data via GRTS for approximately 6 days
between February 2 and February 8 while maneuvers are performed to
stop the TDRS-1 drift. Normal GRO data acquisition via the TDRS
East and West spacecraft will continue during that time.
The performance of the batteries continues to be monitored very
closely, with power configurations adjusted as necessary. The
observatory is functioning normally with primary systems powered
from modular power subsystem 2, which is performing very well.
Modular power subsystem 1 continues to support a reduced load of
mainly heaters.
Preparations are being made to perform a refill of the EGRET
instrument spark chamber gas. This will be the third on-orbit gas
refill for EGRET.
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 01/24/94 through 01/30/94, 2 real time passes were
supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 01/24/94 through 01/30/94, 7 real time passes
were supported with 30 commands transmitted.
A Geospace Environment Modeling (GEM) campaign started on January
11 at 1600Z and ended on January 14 at 1600Z. Data recovered from
IMP-8 during this period was 52.22 percent of requested.
Unfortunately the spacecraft was in the southern hemisphere for the
majority of the time and the University of Tasmania did not support
during any of the time requested.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
Shadow season #33 continues to progress without incident.
A general strike in Spain occurred on January 27. Backup support
for that day's eclipse had been scheduled through Santiago, however
the strike did not affect operations.
On January 27, a data block 17 uplinked to the spacecraft was not
properly processed. The data block was received by the spacecraft
command decoders however it was not completely transferred into the
onboard computer's memory. This type anomaly has been documented
several times over the mission.
A Delta-V is planned for February 15. FDF is still working on
the final details for this orbit adjustment maneuver.
OPERATIONS
A total of 83 images were taken during the week:
Long Wavelength Primary (LWP). . . . . . . . . . 44
Long Wavelength Redundant (LWR). . . . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . . . 39
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minute
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
Steve A.
[email protected]
|
719.6 | February 10-16 | VERGA::KLAES | Quo vadimus? | Wed Feb 16 1994 13:56 | 219 |
| From: US1RMC::"[email protected]" "Steve Andrews" 16-FEB-1994
To: [email protected]
CC:
Subj: Orbiting Satellites
Goddard Weekly
From Code 602, Orbiting Satellites Project
Project Manager Paul J. Pashby
Week of Feb 10, 1994 to Feb 16, 1994
EUVE, ICE, IMP, IUE,
ERBS, COBE, SAMPEX, GRO
EXTREME ULTRAVIOLET EXPLORER (EUVE)
EUVE mission operations proceeded nominally during the reporting
week. Both the Explorer Platform and the EUVE science payload are
operating very well. Observations of the Vela nova remnant and PG-
1520 (ongoing) will be followed by pointings to Capella and TZ-CrB.
SPACECRAFT STATUS
The EUVE spacecraft bus continues to perform very well. All
telemetry values have been within their nominal ranges. A slow
warming of the HGA gimbal Steering Control Electronics has been
observed as expected this week due to the change in HGA management
strategy adopted during the ongoing Shuttle mission. No adverse
consequences of the warming are foreseen. Battery performance
remains very stable. GSFC power engineers suggested in a recent
battery working group meeting that EUVE's transition to constant
current mode at the beginning of orbit day is now largely
unnecessary because of the effective solar array offset angle (40
degrees) kept with the sun. Plans are being reviewed for a return
to normal V/T charging at orbit dawn to be implemented in early
March as an ongoing test.
INSTRUMENT STATUS
No new status or conditions to report.
UPCOMING EVENTS
EUVE will finish the observation of PG-1520 on Feb. 15th, to be
followed by slews to the stars Capella, and TZ-CrB.
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 02/07/94 through 02/13/94, 2 real time passes were
supported with no commands transmitted. One pass on February 9
with station D14 was a demonstration pass to exercise a new
Telemetry Group Controller (TGC-4). Data were degraded during the
entire pass. The station was using a single downlink during this
pass due to a red Maser on the other link.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 02/02/94 through 02/13/94, 7 real time passes
were supported with 29 commands transmitted.
ICE/IMP TRANSP. PROJECT OPS. CONTROL CENTER (TPOCC)
Testing continued by the FOT continued on release 7 throughout this
week. The ATSC test team has completed their testing of release 7.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
Shadow season #33 concluded on February 8. All shadow related
operations were performed without incident during this shadow season.
On February 9 several power hits resulted from the adverse weather
conditions. These power hits had a minimal impact on operations.
During the VILSPA shift on February 10, the VILSPA ground system
was inoperable from 0706 z through 0748 z. At the time their
ground system failed, one of the cameras onboard was in a
configuration which would be undesirable if left for a prolonged
time period. Therefore, the IUEOCC operations director executed a
procedure to place the camera in a safe mode until the VILSPA
station could be made operable again. Although the IUEOCC is
always prepared to take over the command function if VILSPA is
unable to operate, in this instance no further actions by the
IUEOCC staff was required.
On February 11 an instance of the "simultaneous manipulation" of
Worker 18's scheduling counter occurred. When this situation
occurs, Worker 18 (code to process uplinked data blocks) is
incorrectly scheduled to execute its code and the data block
uplinked is not processed properly. This situation results from a
known side-effect in the onboard computer code. In this case, the
situation was immediately identified and action was taken
to resolve the problem. No impact to science operations resulted
from this event.
A 128 count Delta-V was successfully performed February 15 at 11:01
A.M. (local time).
OPERATIONS
A total of 87 images were taken during the week:
Long Wavelength Primary (LWP). . . . . . . . . . 41
Long Wavelength Redundant (LWR). . . . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . . . 46
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minute
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
OPERATIONAL EFFECTIVENESS
Total Time Available : 6720 minutes
Minus losses due to:
Engineering Tests 0 minutes
Spacecraft Problems 0 minutes
Ground System Problems 0 minutes
Operations Problems 0 minutes
Time used in Science Data Ops: 6720 minutes
Effectiveness 100%
PROBLEMS AND CONCERNS
Recent attempts to use the ground simulator for the onboard
computer have failed. Maintenance has given a preliminary
diagnosis that the memory has failed in the unit. This will
severely impact the testing of any software modifications needed to
the onboard computer. At this time no software modifications are
anticipated; however, if an unexpected need for a software patch
arises there would be no means of functionally testing
the modification.
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
has remained oriented toward the A0535+26 target of opportunity
since February 8. The observation of this target was extended
beyond the expected one week because the source has continued to
brighten and interesting structure in the spectrum is being
detected. An adjustment to the attitude was performed on February
15 to improve the sun angle on the arrays. The next attitude
maneuver is planned for February 17 to the NGC 1068 science target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of
full-time coverage achieved with 32 Kbps real-time data from
February 2 through February 8 was 76 percent. An additional 5
percent coverage was achieved with the BATSE 1 Kbps subset of data
during this time. The acquisition of GRO data via the GRO Remote
Terminal System (GRTS) was suspended for approximately 6 days
between February 2 and February 8 while maneuvers were performed to
stop the TDRS-1 drift. TDRS-1 is now in the vicinity of its
on-station location of 85 degrees East longitude, and GRO support
via GRTS has resumed while engineering evaluation of GRTS continues.
The performance of the batteries continues to be monitored very
closely, with power configurations adjusted as necessary. The
observatory is functioning normally with primary systems powered
from modular power subsystem 2, which is performing very well.
Modular power subsystem 1 continues to support a reduced load of
mainly heaters.
The third on-orbit refill of the EGRET instrument spark chamber gas
was successfully performed February 8-9.
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft is still operating with only one battery with
two failed cells, while acquiring scientific data from both the
ERBE-NS instrument and the SAGE instruments. The spacecraft began
to experience orbital eclipses again last week after 10-days of
continuous sunlight. A spacecraft yaw-turn is scheduled for late
next week.
COSMIC BACKGROUND EXPLORER (COBE)
The COBE Mission was terminated on January 28th, and the spacecraft
has been transferred to TOTS Project at Wallops and to the STGT
Project here at GSFC for them to use as a communications test
satellite for their ground stations. Both Projects have
successfully commanded the spacecraft On & Off repeatedly, and
the COBE spacecraft has been used as a test satellite over a
scientific ground station in Antarctica for Code 450. [TOMS Project]
SAMPEX
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. One new
Front-End-Processor was installed in the Control Center, along
with new SAMPEX software one week ago. The Flight Ops Team is
currently testing this new software in a parallel with the ongoing
SAMPEX operations. Two more Front-End-Processors are expected to
be installed by early March, and the new FAST software will be
integrated into the Control Center soon afterwards. The FAST space-
craft is scheduled to be launched in six months.
[email protected]
|
719.7 | February 17-23 | VERGA::KLAES | Quo vadimus? | Fri Feb 25 1994 12:12 | 176 |
| From: US1RMC::"[email protected]" "Steve Andrews" 25-FEB-1994
To: [email protected]
CC:
Subj: Orbiting Satellites
Orbiting Satellites Project Code 602
Project Manager: Paul J. Pashby
Week of: Feb 17, 1994 to Feb 23, 1994
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 02/14/94 through 02/20/94, 2 real time passes were
supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 02/14/94 through 02/20/94, 7 real time passes
were supported with 20 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
The patches to the onboard computer code required for the proper
operation during shadow seasons were removed on February 14.
A 128 count (0.1024 seconds/count) orbit adjustment maneuver was
performed on February 15. Of the 13.1072 second total burn time,
1.28 seconds was lost due to the spacecraft body angles exceeding
the control switching line limits. The preparation prior to the
actual maneuver along with the attitude recovery process required
following the maneuver required a total time of three hours seven
minutes (187 minutes). Following attitude recovery, science
operations resumed as normal.
On February 16, the building maintenance engineers tested the
diesel power system. Prior preparation for this test by the OCC
staff prevented this test from significantly affecting science
operations. Equipment essential for operations was switched from
the load center being tested to another load center.
On February 18 extreme interference problems with the telemetry
downlink signal were experienced between 1840 Z and 1926 Z. The
interference originated from Naval equipment transmitting at almost
the same frequency as IUE. The IUE telemetry system was transferred
from the SBAND to the VHF system in order to reacquire the downlink
signal, allowing operations to continue without interruption.
The weekly report covering February 7 through February 13 reported
that the simulator for the onboard computer was not operational.
The malfunction in this equipment has been corrected and the
simulator is operational at this time.
OPERATIONS
A total of 98 images were taken during the week:
Long Wavelength Primary (LWP). . . . . . . . . . 39
Long Wavelength Redundant (LWR). . . . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . . . 59
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . .187 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minute
Operations Problems. . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6593 minutes
Effectiveness. . . . . . . . . . . . . . . .97.2%
EXTREME ULTRAVIOLET EXPLORER (EUVE)
The Extreme Ultraviolet Explorer continues to perform well. A
single event upset was experienced on the payload Central Command,
Data, and Power Unit (CDP), but it had no effect on the ongoing
observation. EUVE will conduct observations of WD-1057+719 and
AG-Dra in the coming week.
SPACECRAFT STATUS
The EUVE Explorer Platform continued its overall good performance.
Intermittent guide star acquisition problems in star tracker #2
have been experienced since slewing to WD-1057+719 on the evening
of the 20th, but this problem should not significantly compromise
the quality of the observation. Similar problems occurred on Jan.
6-7 when this target was also observed and are most likely
attributable to numerous additional stars in the tracker's field of
view. Thermistor #3 on solar array #1 was also reporting elevated
temperatures at this spacecraft attitude - this caused by specular
reflection from the spacecraft body. Ground-commanded reorien-
tation commands to the arrays promptly lowered the temperature.
This incident will be studied closely, as the temperature reported
on solar array #1 exceeded the predicted temperature at that pitch,
roll and solar array offset.
INSTRUMENT STATUS
The daily memory dump of the payload from orbit 9436 (Feb. 18)
indicated that the CDP software experienced an SEU. During orbit
9440 the CDP was successfully reloaded. There was no effect on the
observation and no science data lost.
UPCOMING EVENTS
EUVE will finish observing WD_1057+719 and then slew to AG_Dra
later in the week.
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the NGC 1068 science target on February 17.
The next attitude maneuver is planned for March 1 to the QSO
0716+714 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of
full-time coverage achieved with 32 Kbps real-time data from
February 9 through February 15 was 80 percent, including 18 percent
from the GRO Remote Terminal System (GRTS). An additional 1
percent coverage was achieved with the BATSE 1 Kbps subset of data
during this time. Engineering evaluation of GRTS continues.
The performance of the batteries continues to be monitored very
closely, with power configurations adjusted as necessary. The
observatory is functioning normally with primary systems powered
from modular power subsystem 2, which is performing very well.
Modular power subsystem 1 continues to support a reduced load of
mainly heaters.
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft is still operating with only one battery with
two failed cells, while acquiring scientific data from the ERBE-NS
instrument most of the time and from the SAGE instrument all of the
time. The spacecraft will be yawed 180 degrees later this week in
order to keep sunlight on the solar arrays as the sun moves to the
other side of the spacecraft. The ERBE-NS instrument is currently
powered OFF for about one week during the longest eclipses.
COSMIC BACKGROUND EXPLORER (COBE)
The COBE Mission was terminated on January 28th, and the spacecraft
has been transferred to TOTS Project at Wallops and to the STGT
Project here at GSFC.
SAMPEX
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. A leak test
was performed on the HLT instrument this past week. Also several
minor anomalies occurred on the spacecraft this past week. Several
noise pikes occurred on the high votlage in the Leica instrument,
and a "warm restart" occurred in the spacecraft telemetry system.
The Control Center is still undergoing changes in preparation for
the FAST launch in August. Another Front-End-Processor will be
upgraded next week, and the new SAMPEX software is still being
tested in parallel to the normal SAMPEX operations.
Steve A.
[email protected]
|
719.8 | March 10-16 | JVERNE::KLAES | Be Here Now | Wed Mar 23 1994 12:05 | 136 |
| Article: 5442
From: Steve Andrews <[email protected]>
Newsgroups: sci.space.news
Subject: Orbiting Satellites 3/10/94
Date: 23 Mar 1994 06:38:53 -0800
Organization: NASA Goddard Space Flight Center
Sender: [email protected]
Orbiting Satellites Project Code 602
Project Manager Paul J. Pashby
From March 10, 1994 to March 16, 1994
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
will be oriented toward the QSO 0716+714 science target on March 15. The
next attitude maneuver is planned for March 22 to the GAL 357-11 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of full-time
coverage achieved with 32 Kbps real-time data from March 2 through March 8
was 80 percent, including 18 percent from the GRO Remote Terminal System
(GRTS). An additional 2 percent coverage was achieved with the BATSE
1 Kbps subset of data during this time. Engineering evaluation of GRTS
continues.
The EGRET instrument was switched from the MPS-2 power source to
MPS-1 on March 8 to provide better power load management for both
MPS-1 and MPS-2. At the same time, one of the two reaction wheels
that was placed on MPS-1 for orbit reboost was returned to MPS-2.
Battery operation is stable.
Beginning on March 13, a subset of the Comptel instrument
engineering telemetry points began to indicate they were out of
limits; however, the science data stream telemetry verified the
instrument was operating properly. The instrument remains in a
normal operating condition while engineering analysis is being
performed to isolate the cause of the apparently erroneous
telemetry readings.
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 03/07/94 through 03/13/94, 2 real time passes were
supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 03/07/94 through 03/13/94, 7 real time passes
were supported with 31 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
On March 7, line problems between the Wallops tracking station and
the IUEOCC occurred lasting several minutes; however, these
problems did not impact the operations in progress at the time.
On March 10, problems with the Sigma 5 ground system computer
resulted in a loss of 50 minutes of science time. The Sigma 5
system crashed and would not reboot, forcing operations to be
switched to the backup Sigma 9 ground system computer.
OPERATIONS
A total of 111 images were taken during the week:
Long Wavelength Primary (LWP). . . . . . . . . . 80
Long Wavelength Redundant (LWR). . . . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . . . 31
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering Tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . 50 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops:. . . . . . 6670 minutes
Effectiveness. . . . . . . . . . . . . . . .99.3%
EXTREME ULTRAVIOLET EXPLORER (EUVE)
EUVE continues to operate well. No significant problems or
incidents to report. The spacecraft will conduct routine Guest
observations until the end of the week when it will devote two days
to the collection of engineering trend and calibration data- this
to be done in both Initial Point and Survey Modes.
SPACECRAFT STATUS
The spacecraft Explorer Platform performed nominally this week.
Minor changes in several battery parameters - including differential
voltage signatures, charge current profiles, and the end of night
load bus voltage - prompted GSFC Flight Controllers to alter the
battery management strategy for an orbit to allow more charging.
A fine balance is continually sought between adequate charge and
minimal overcharge. Initial battery response was good; a battery
working group meeting this week will discuss future plans.
Overall, battery performance remains very good.
The spacecraft will transition to Survey Mode at the end of the
week to collect quarterly trend data on the performance of the star
trackers. This will be followed by approximately twenty spacecraft
slews in Inertial Mode which have been carefully chosen to move the
sun through regions of the Digital Fine Sun Sensor which do not
routinely see the sun. This effort is being done to characterize
the sensor's performance across its entire field of view while the
mission's two star trackers are performing well and can be used for
reference. This is a proactive measure; there is nothing in the
star tracker performance to indicate imminent failure.
INSTRUMENT STATUS
An "allwsz" test was conducted for three orbits this week to study
the effect on the detectors. Commands to normalize the
configuration at the test's end were accomplished successfully.
UPCOMING EVENTS
EUVE will transition to Survey Mode briefly at week's end- followed
by a short set of engineering slews in Inertial Mode. Science
observations will recommence at approximately 0830Z on the 22nd.
Steve A.
[email protected]
|
719.9 | March 17-23 | JVERNE::KLAES | Be Here Now | Wed Mar 23 1994 16:02 | 180 |
| Article: 5443
From: Steve Andrews <[email protected]>
Newsgroups: sci.space.news
Subject: Orbiting Satellites 3/17/94
Date: 23 Mar 1994 11:07:02 -0800
Organization: NASA Goddard Space Flight Center
Sender: [email protected]
Orbiting Satellites Project Code 602
Project Manager Paul J. Pashby
From March 17, 1994 to March 23, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE)
The EUVE spacecraft and its operation both continue to go well.
Guest observations of the bright star Procyon and the Vela Supernova
Remnant were followed by a transition to Survey Mode for the
purpose of gathering engineering trend data on the fixed head star
trackers. At the conclusion of the engineering time, which also
includes time for calibration of the Digital Fine Sun Sensor
(DFSS), EUVE will conduct an observation of the planet Jupiter.
SPACECRAFT STATUS
The spacecraft Explorer Platform performed nominally this week-
telemetry values from all subsystems were in the expected ranges.
Much planning was completed for the engineering test period which
the spacecraft entered at the end of the week. After gathering
star tracker performance data in Survey Mode, the spacecraft will
transition back to Inertial Mode and execute 20 maneuvers designed
to sweep the sun across the field of view of the DFSS. Analysis of
the data will be accomplished by the GSFC Flight Dynamics Facility
with the goal of relating the sun sensor data gathered to the
reference frame defined by the star trackers. At the end of the
engineering test time, three software flags in S/C table 88 will be
disabled. This will relax the attitude control of the spacecraft,
allowing normal gravity and drag-related orbital forces to gently
nudge the S/C from its target vector. This test is being done in
cooperation with Univ. of California Berkeley personnel as a
possible low-effort means to average out the fixed pattern noise of
the payload spectrograph. A successful result could eliminate the
need for multiple pointings (dithers) currently done on candidate
targets. Recognition is given to the Flight Ops. Team, FDF, and
Fairchild Space personnel for their careful planning and execution
of the entire two day engineering test period.
INSTRUMENT STATUS
No new status or incidents reported by UCB.
UPCOMING EVENTS
At the conclusion of the engineering test period in the early hours
of the 22nd, EUVE will slew to Jupiter to begin a three day observation.
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GAL 357-11 science target on March 22. The
next attitude maneuver is planned for April 5 to the MRK 421 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of
full-time coverage achieved with 32 Kbps real-time data from March
9 through March 15 was 77 percent, including 15 percent from the
GRO Remote Terminal System (GRTS). An additional 3 percent coverage
was achieved with the BATSE 1 Kbps subset of data during this time.
Engineering evaluation of GRTS continues. GRTS is expected to be
declared fully operational on April 1, 1994.
The erroneous subset of Comptel instrument engineering telemetry
parameters that began on March 13 appears to have been caused by an
internal circuit failure within Comptel's Remote Interface Unit B
(RIU-B). On March 17, the redundant RIU-A was switched on-line,
and all telemetry measurements returned to normal. The observatory
will remain in this configuration. During the time of the erroneous
readings, the Comptel instrument continued to operate normally and
there was no impact to science data.
Preparations are being made for a transition of the GRO Flight
Operations Team from TRW to the NMOS/Allied Signal contract. The
formal 3-month phase-in will begin April 7, 1994. Contract transition
will occur on July 7, 1994.
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis. During
the week of 03/14/94 through 03/20/94, 2 real time passes were
supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 03/14/94 through 03/20/94, 7 real time passes
were supported with 30 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
On March 14 the spacecraft experienced an anomaly that has
occurred several times in the past. An exposure controlled by the
Onboard Computer was terminated by the computer 2 "ticks" (.8192
seconds) early. As the total exposure time was 333.3 minutes, the
lost ticks did not effect the image, and science operations were
not impacted. This was recorded as Spacecraft Anomaly G-198.
On March 17, handover to the VILSPA station was delayed by 3
hours and 35 minutes due to equipment problems at the VILSPA
station. As VILSPA could not command, personnel remained in our
SOC and executed the first maneuver of VILSPAs planned program.
The ISLU backup commanding & telemetry device was used from
0510z to 0635z so VILSPA could command the spacecraft through our
SIGMA computer. At 0635z VILSPAs equipment was put on line and
normal operations resumed.
On March 19, due to a scheduled power outage at Wallops, we ran
operations through the Santiago (AGO) backup station from 1230z
until 2155z. There was a 10 minute loss of telemetry which was
resolved when a switch to normal telemetry from convoluted
telemetry was made. Also, due to various problems, 129 minutes of
science time were lost. The problems were mostly due to the
Spacecraft Commanding Equipment (SCE). The SCE had to be reloaded
at one point, and in one instance sent a block of commands more
than once. Although the spacecraft was in no danger it did cause
a temporary loss of attitude reference (the double commands were
for a very small maneuver).
OPERATIONS
A total of 137 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 63
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . .. 74
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . .129 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6591 minutes
Effectiveness. . . . . . . . . . . . . . . .98.1%
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft is still operating with only one battery with two
failed cells, while acquiring scientific data from both the ERBE-NS
and SAGE instruments. The Flight Ops Team are modifying the mission
planning software for changes that were requested by the LaRC
experimenters. IPD (Code 562) is now sending the ERBS Ephemeris data
to LaRC electronically. The Weekly Ephemeris tapes will be terminated
in a few weeks if everyone is happy with the new method.
SAMPEX
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The Control
Center is still being upgraded for the new FAST mission which is
scheduled to be launched in August. Two days of simultaneous SAMPEX
and FAST operations were successfully run in the Control Center
this week. The ACS engineers (Codes 712.3) and the Flight Software
engineers (512.2) are still testing the new ACS pointing algorithms
for the SAMPEX spacecraft which were requested by the Project
Scientist. The HLT instrument operations have been changed to start
conserving the isobutane gas for the possible 3-yr extension to the
SAMPEX mission.
Steve A.
[email protected]
|
719.10 | March 24-31 | GLITTR::KLAES | Be Here Now | Sun Apr 03 1994 16:51 | 185 |
| Article: 5484
From: Steve Andrews <[email protected]>
Newsgroups: sci.space.news
Subject: Orbiting Satellites 3/24/1994
Date: 1 Apr 1994 06:25:03 -0800
Organization: NASA Goddard Space Flight Center
Sender: [email protected]
CODE 602 Orbiting Satellites Project
Prokect Manager: Paul J. Pashby
Week of: March 24, 1994 to March 31, 1994
COMPTON GAMMA RAY OBSERVATORY (GRO)
The Compton observatory is in normal pointing mode. The +Z axis
has been oriented toward the GAL 357-11 science target since
March 22. The next attitude maneuver is planned for April 5 to
the MRK 421 target.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from March 16 through March 22 was 80 percent, including 17
percent from the GRO Remote Terminal System (GRTS). An
additional 5 percent coverage was achieved with the BATSE 1 Kbps
subset of data during this time. Engineering evaluation of GRTS
continues. Although there will still be periods of no user
support reserved for maintenance and software deliveries, GRTS is
expected to be declared fully operational on April 1, 1994.
Preparations continue for a transition of the GRO Flight
Operations Team from TRW to the NMOS/AlliedSignal contract. The
formal 3-month phase-in will begin April 7, 1994. Contract
transition will occur on July 7, 1994.
Revision 3 of the GRO Support Instrumentation Requirements
Document has been signed by NASA Headquarters Codes O and S.
EARTH RADIATION BUDGET SATELLITE (ERBS)
The ERBS spacecraft continues to operate with only one battery
with two failed cells. The SAGE Instrument continues to acquire
science data at every orbital sunset & sunrise. The ERBE-NS
instrument will be powered OFF for the next 7-days in order to
reduce power during the time of maximum eclipses. The spacecraft
will also be yawed 180 degrees later this week in order to keep
sunlight on the solar arrays as the ERBS orbit precesses with
respect to the sun.
COSMIC BACKGROUND EXPLORER (COBE)
The Last 6-Month Trending Report was distributed last week, and
the End-of-Life Test Report will be released this week. All of
the COBE documentation & flight records have been cleared out,
and the three remaining Control Center personnel will be
reassigned to other Projects later this week.
SAMPEX
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
Control Center is still being upgraded for the new FAST mission
which is scheduled to be launched in August. The HILT instrument
has been cycled ON/OFF several times in the last few weeks in
order to check how fast the isobutane gas leaks out of the
instrument. The experimenter is attempting to conserve the
isobutane gas for the extended SAMPEX mission. The instrument
will continue to acquire data during the 3-month cycles that the
isobutane is turned OFF.
EXTREME ULTRAVIOLET EXPLORER (EUVE)
The EUVE spacecraft marked its 10,000th revolution on-orbit this
past week by slewing to observe a bright, unexpected Target of
Opportunity in the direction of the brilliant B-type star Hadar
(Beta Cen) in the southern constellation of Centaurus. All
payload and platform subsystems continue to work flawlessly. Two
days will be spent viewing this surprising source; EUVE will then
rejoin the prescheduled science plan with observations of HZ 43
and Alpha Centauri.
SPACECRAFT STATUS
Performance of the EUVE Explorer Platform continues to be very
good. The engineering test period conducted at the beginning of
the week went exceedingly well. Smooth transitions were conducted
in and out of Survey Mode to gather performance trend data on the
fixed head star trackers. Twenty special slews which followed
(designed to calibrate the field of view edges of the fine sun
sensor) also went without incident. The data collected during
the relaxed attitude control test (which concluded the
engineering period) is now being assessed. No unexpected
spacecraft movement or software responses occurred, but more
tests are likely to validate this possible alternate method of
dithering.
INSTRUMENT STATUS
On the 26th, an unexpectedly bright target was observed in
scanners 1 and 2. Scanner #3 had been turned off to increase the
telemetry throughput for the Deep Survey/Spectrometer which was
viewing the target HZ 43. Scanner #3 was also turned on to
observe the bright source and after analysis and consultation
with the GSFC Project Scientist Office, it was decided to treat
the source as a target of opportunity and reorient the spacecraft
to allow its observation by the DSS. Analysis of the spectral
data is now ongoing.
UPCOMING EVENTS
Observations of Alpha Centauri and J1104+38.2 are scheduled at
the conclusion of the present exposure on the unidentified target
of opportunity.
INTERNATIONAL COMETARY EXPLORER (ICE)
Support of the ICE spacecraft continues on a routine basis.
During the week of 03/21/94 through 03/27/94, 2 real time passes
were supported with 2 commands transmitted.
The Steinberg experiment (SBH) on the ICE spacecraft was turned
off, on March 21,1994, to conserve power as the spacecraft moves
further than 1AU from the Sun.
INTERPLANETARY MONITORING PLATFORM (IMP)
Support of the IMP-8 spacecraft continues on a routine basis.
During the week of 03/21/94 through 03/27/94, 7 real time passes
were supported with 25 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE)
HIGHLIGHTS
On March 22, due to an unavailable 18m antenna at Wallops and the
spacecraft pointing angle, the telemetry signal was not strong
enough to support camera reads. This caused a loss of 8 minutes
of science time.
On March 25 a possible spacecraft anomaly occurred when a SCAN
command to one of the cameras did not take effect. This is
currently under investigation. The command was sent again and
worked correctly. Normal operations resumed within 5 minutes.
On March 26 two possible spacecraft anomalies occurred and are
under investigation. First, a command to switch S-BAND antennas
did not take effect, and secondly, a command to slew the
spacecraft did not take effect. Both commands were retransmitted
and worked correctly. Operations were not effected.
OPERATIONS
A total of 160 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 33
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . ..127
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .5 minutes
Ground System Problems . . . . . . . . .8 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6707 minutes
Effectiveness. . . . . . . . . . . . . . . .99.8%
Steve A.
[email protected]
|
719.11 | This just is not right | JVERNE::KLAES | Be Here Now | Mon Apr 04 1994 14:12 | 19 |
| Article: 55767
From: [email protected] (Richard A. Schumacher)
Newsgroups: sci.astro
Subject: Turning off active satellites (Was Re: Mars Observer II?)
Date: 3 Apr 94 21:21:42 GMT
Organization: CONVEX Computer Corporation, Richardson, TX USA
The May 1994 issue of S&T has a letter from Jeffrey Linsky of the
IUE User's Committee on this subject. Under present NASA policy the
International Ultraviolet Explorer, Rosat, Extreme UV Explorer and the
Compton Gamma Ray Observatory would all be turned off soon after
September 1994, ostensibly to save money.
Meanwhile, space station activities to date have spent eleven billion
dollars with no science return. Another eighteen billion dollars, more
or less, are planned to be spent. Let your representatives know that
you care about getting our money's worth from our investment in
science missions.
|
719.12 | May 5-12 | MTWAIN::KLAES | Keep Looking Up | Mon May 23 1994 12:41 | 312 |
| From: US4RMC::"[email protected]" "Steve Andrews" 23-MAY-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 5/5/1994
CODE :602
INDIVIDUAL :Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF : May 5, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
The EUVE observatory continues to prove itself robust with the
successful completion of two days of engineering tests conducted
outside its normal pointing constraints. All payload and
platform subsystems are performing well. Another attempt will be
made this week to detect any residual RF output from the failed
transmitter "B" using EUVE's omni antennas and the MILA tracking
station. Science targets for the coming week include HR-4804 and
PKS-2155.
SPACECRAFT STATUS
Two days of special engineering tests were conducted May 7-9 to
determine the spacecraft's ability to perform safely at attitudes
outside the mission's normal pointing constraints. Maneuvering to
these attitudes will be necessary to observe the planet Jupiter in the
early August time period- just following the collision of Comet
Shoemaker-Levy with that planet. Much careful planning went into the
series of eight slews which, in their course, exercised the spacecraft
at various pitch and roll angles previously unexperienced on orbit.
Ground-commanded repositioning of the solar arrays, as well as the
removal of constant current control mode at the beginning of orbit day
was necessary at some points to keep the observatory power positive.
Further analysis of the engineering data is underway, but initial
results indicate no negative performance on the part of the Explorer
Platform or the science payload. There were some periods of guide
star acquisition problems in both star trackers, but these are
believed to result from the presence of numerous blocker stars in the
fields of view of the trackers. Congratulations to the members of the
Flight Operations Team for their careful planning and execution of
these tests, and thanks to the UCB SOC and GSFC Flight Dynamics Group
and personnel from Fairchild Space for their support and assistance.
In other matters, a charging efficiency value in spacecraft Table 28
was changed on 5/2 from .98 to .97 to increase the end of night load
bus voltage and decrease the magnitude of differential voltage
"tail-ups" recently seen at the end of night. This is just another
step in the careful management of the batteries- no serious battery
problems are indicated. Two more tests with the MILA tracking station
are being planned for May 11 to listen for any low level RF output
from EUVE's transmitter "B" which failed on April 2nd. Permission has
been granted by the Associate Project Scientist to actually reposition
the spacecraft into the best geometry for the ground test during the
second of these opportunities. This should all be accomplished in the
daylight portion of the orbit, so no science data loss is anticipated.
INSTRUMENT STATUS
No new incidents or problems to report. Tests to study the effect of
"wedge-strip-zigzag" (wsz) high voltage increases on the detector gain
are ongoing. Analysis of the DSS performance during the engineer test
is also underway.
UPCOMING EVENTS
Two ground tests will be conducted with MILA on the 11th. Guest
Observer targets HR-4804 and PKS-2155 will be observed this week.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis was
oriented toward the NGC 3227 science target on May 10. The next
attitude maneuver is planned for May 17 to the GAL 083-45 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of full-time
coverage achieved with 32 Kbps real-time data from April 27 through
May 2 was 93 percent, including 9 percent from the GRO Remote Terminal
System (GRTS). The high- gain antenna boom angle with respect to the
equatorial plane was 80 degrees.
The phase-in period of the GRO Flight Operations Team from TRW to
the NMOS/AlliedSignal contract is proceeding normally.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to operate with only one battery with
two failed cells, while acquiring science data from both the SAGE and
ERBE-NS instruments. NASA Hqtrs held an ERBS Project Review at GSFC
this week. They reviewed the spacecraft and instrument status, the
mission operations, the data processing, the scientific achievements,
and the ERBS budget. They agreed that ERBS data is valuable and the
spacecraft should be kept operating as long as possible.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data without
any significant problems or operational errors. The tests for the
change to the onboard attitude control software have been completed by
the Flight Systems Branch (Code 512), and the test results and the
proposed patches have been reviewed by SAMPEX Flight Software CCB.
However, a few more tests to evaluate the effect of "warm restarts"
will be performed this week in order to help alleviate some concerns
of the Flight Software Branch (Code 735.3). The SMEX Control Center
supported several FAST mission simulations this week as a part of the
ongoing FAST Thermal Vacuum tests. The High Voltage problem on the
Leica instrument reoccurred about six times this week. There are
safety procedures that are automatically invoked onboard the
spacecraft when this condition occurs.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis. During the
week of 05/02/94 through 05/08/94, 2 real time passes were supported
with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine basis. During
the week of 05/02/94 through 05/08/94, 7 real time passes were
supported with 42 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (ICE) - G. Bullock
HIGHLIGHTS
On May 2, 1994, fifty minutes of telemetry were lost due to equipment
problems at the OCC. This outage did not impact science operations.
On May 3, 1994, Five minutes of telemetry were lost due to equipment
problems at the OCC. Also, on May 4, a total of 9 minutes were lost
due to equipment problems at both Wallops and the OCC. All of these
losses effected science operations.
OPERATIONS
A total of 113 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 60
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 53
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . .6720 minutes
Minus losses due to:
Engineering Tests . . . . . . . . . . 0 minutes
Spacecraft Problems . . . . . . . . . 0 minutes
Ground System Problems. . . . . . . . 9 minutes
Operations Problems . . . . . . . . . 0 minutes
Natural Phenomena . . . . . . . . . . 0 minutes
Time used in Science Data Ops: 6711 minutes = 99.87 %
Steve A.
[email protected]
From: US4RMC::"[email protected]" "Steve Andrews" 23-MAY-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 5/12/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF : May 12, 1994
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GAL 083-45 science target on May 17. The
next attitude maneuver is planned for May 24 to the PSR 1951+32 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of full-time
coverage achieved with 32 Kbps real-time data from May 3 through May 9
was 93 percent, including 9 percent from the GRO Remote Terminal
System (GRTS). The high- gain antenna boom angle with respect to the
equatorial plane was 80 degrees.
The phase-in period of the GRO Flight Operations Team from TRW to
the NMOS/AlliedSignal contract is proceeding normally.
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
EUVE conducted a week of nominal operations on targets HR_4804
and PKS-2155. All subsystems are operating well. Two attempts
to acquire any weak residual signal from transmitter "B" using
the MILA tracking station were both unsuccessful. Targets for
the coming week include the hot white dwarf star WD1314.
SPACECRAFT STATUS
Two tests with the MILA tracking station were conducted on May 11
to listen for any low level RF output from EUVE's transmitter "B"
which might still be detectable when the transmitter is turned
on. Care was taken to "maximize the geometry" between the
spacecraft and the ground so as to minimize the presence of back
or sidelobe radiation from the high gain antenna. As a result,
periods of both passes- which were also both high elevation
opportunities over MILA- were effectively free from the presence
of a TDRSS mode signal from the high gain. Unfortunately,
neither attempt acquired any GSTN carrier. This means the
signal- if there at all- is below a value of -31dBm which implies
at least a 60 dB transmitter loss. Further analysis of the data
will be undertaken before a decision is made to schedule any more
ground tests.
A telecon was held with UCB during the week to coordinate plans
for the testing of UCB-supplied Absolute Time Commands. Two
tests with the GSFC SDVF simulator are planned for the coming
week. Tests between UCB and the GSFC CMS systems continue in
parallel. Work continues as well between UCB and GSFC on final
plans for the Jupiter/SL9 observations later in the summer.
INSTRUMENT STATUS
No new incidents or problems to report. Tests to study the
effect of "wedge-strip-zigzag" (wsz) high voltage increases on
the detector gain were recently concluded, while analysis of the
data taken is underway.
UPCOMING EVENTS
Guest Observer target WD1314 is scheduled to be observed this week.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis.
During the week of 05/09/94 through 05/15,94, real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 05/09/94 through 05/15/94, 7 real
time passes were supported with 29 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
On May 9, a total of 49 minutes of telemetry was lost due
to ground equipment problems at both the IUEOCC and at
Wallops. About 30 minutes of this lost time occurred when
the THS on the Sigma 5 failed, and a switch to the Sigma 9
was needed. The remaining time loss was due to telemetry
line problems between OCC and WPS. Approximately 30 of the
49 minutes actually impacted science operations.
On May 13, a 20 minute exposure was lost when an
unexplained telemetry loss occurred during its downlink.
The downlink signal was excellent, however WPS insisted that
the telemetry hits came from the spacecraft.
On May, the 18m telemetry antenna went down at Wallops
and it was 11 min before a backup 7.5m antenna could be put on line.
OPERATIONS
A total of 91 images was taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 42
Long Wavelength Redundant (LWR) . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 49
OPERATIONAL EFFECTIVENESS
Total Time Available :. . . . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . 61 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6659 minutes
Effectiveness. . . . . . . . . . . . . . . 99.09%
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to operate with only one battery
with two failed cells, while acquiring science data from both
the SAGE and ERBE-NS instruments. The spacecraft will begin a
12-day period of continuous sunlight this week. During these
periods, which occur 4X a year, the SAGE instrument is powered
OFF because no sunrises or sunsets occur on which to acquire
scientific data.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
proposed ACS change to the onboard software is still being tested
on the ground to help alleviate some concerns of the Flight
Software Branch (Code 735.3). This change, which was requested by
the SAMPEX Project Scientist, will greatly increase the science
return from the PET instrument. The HV problem in the LEICA
instrument is still occurring, but the experimenter now feels it
is more due to the increased solar activity than to a further
degradation in the instrument.
Steve A.
[email protected]
|
719.13 | May 19 | MTWAIN::KLAES | Keep Looking Up | Fri May 27 1994 14:11 | 168 |
| From: US4RMC::"[email protected]" "Steve Andrews" 27-MAY-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 5/19/94
CODE : 602
INDIVIDUAL : Paul J. Pashby
TITLE : Orbiting Satellites Project
WEEK OF : May 19, 1994
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine
basis. During the week of 05/16/94 through 05/22/94, 2 real
time passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 05/16/94 through 05/22/94,
7 real time passes were supported with 29 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
On 5/18 the prime command transmitter failed at Wallops
and a backup transmitter had to be configured,
resulting in a 4 minute loss of command capability and
a corresponding loss of science time.
On 5/21 the prime command transmitter failed again at
Wallops and a backup transmitter had to be configured,
resulting in a 8 minute loss of command capability and
science time.
On 5/22 the disc pack on the Sigma 9 failed resulting
in a 3 minute data loss while changing disc packs.
OPERATIONS
A total of 215 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 110
Long Wavelength Redundant (LWR) . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 105
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests . . . . . . . . .0 minutes
Spacecraft Problems . . . . . . . .0 minutes
Ground System Problems. . . . . . 15 minutes
Operations Problems . . . . . . . .0 minutes
Natural Phenomena . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6705 minutes
Effectiveness. . . . . . . . . . . . . . . 99.78%
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z
axis was oriented toward the PSR 1951+32 science target on May
24. The next attitude maneuver is planned for May 31 to the
GAL 253-42 target.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps
real-time data from May 11 through May 16 was 79 percent,
including 17 percent from the GRO Remote Terminal System
(GRTS). The high-gain antenna boom angle with respect to the
equatorial plane was 5 degrees.
Beginning around May 27, limited amounts of GRO support will
be scheduled routinely on the Second TDRS Ground Terminal to
assist in engineering evaluation.
The phase-in period of the GRO Flight Operations Team from TRW
to the NMOS/AlliedSignal contract is proceeding normally
toward a July 7, 1994 transition. Of the 21 incumbent
technical personnel, 14 have accepted positions with
AlliedSignal and two more are pending.
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
The Extreme Ultraviolet Explorer is operating well; no
problems or unusual incidents to report. Targets for the
coming week include WD1314 (HZ43) and RE2127 (EX-HYA).
SPACECRAFT STATUS
Telemetry data from the Explorer Platform's power, C&DH, and
attitude control subsystems was all within the expected ranges
during the reporting week. All systems continue to perform
very well. A lower than usual depth of discharge on the
batteries (11%) occurred on day 140- this the result of the
solar arrays repositioning themselves more quickly than the
spacecraft during a large (almost 180 deg.) spacecraft slew.
No adverse effects on the batteries are expected. Flight
software (FSW) testing was conducted this week between the
POCC and the SDVF spacecraft simulator to check the load and
verification procedures associated with several minor FSW
patches, and to continue the recent progress in testing the
process of receiving absolute time command information from
UCB and merging this with the daily EP ATC loads.
More testing will be done in the coming week.
Planning also continues on the observation of the
Jupiter/comet collision in mid to late July. With the success
of the recent engineering test of roll angles slightly beyond
our normal pointing constraints, observations are now possible
throughout the pre-to-post impact phases at spectrometer
orientations very close to ideal. Moon presence in the DSS
and alternately Star Tracker #1 significantly hinders
observations planned during the period 16-21 July, and work
continues to maximize the observing during this time.
Plans are also being coordinated with the GSFC Flight Dynamics
Group to conduct a series of six special slews on June 3rd to
recalibrate EUVE's gyros and collect quarterly star tracker
trend data in Survey Mode.
INSTRUMENT STATUS
No new incidents or problems to report.
UPCOMING EVENTS
Guest Observer target WD1314 is scheduled to be observed this
week, followed by RE2127. Twenty eight hours are being set
aside on June 3-4 for gyro recalibration and star tracker
trend data collection.
EARTH RADIATION BUDGET SATELLITE (ERBS) - James Williamson
The ERBS spacecraft continues to operate with only one battery
with two failed cells, while acquiring science data from both
the SAGE and ERBE-NS instruments. The spacecraft is currently
in a 12-day period of continuous sunlight. The SAGE instrument
is powered OFF during this period because it would over-heat
and because no orbital sunrise or sunset events occur which
SAGE requires.Special battery charge control procedures are
used in this period to prevent the battery from being over-charged.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
ACS change to the onboard software will be uplinked to the
spacecraft later this week. This software patch has been
thoroughly tested and approved by the SAMPEX Flight Software
CCB. The HLT instrument in now back ON after being OFF for
about one month while the experimenter worked out a procedure
to extend the life of the isobutane gas for the extended phase
of the SAMPEX mission. The SMEX Control Center is still
running many tests in preparation for FAST which is less than
three months to launch.
Steve A.
[email protected]
|
719.14 | June 2-8 | MTWAIN::KLAES | Keep Looking Up | Wed Jun 08 1994 14:40 | 176 |
| From: US4RMC::"[email protected]" "Steve Andrews" 8-JUN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 6/2/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF : June 2 to June 8, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
June 7th marks the second anniversary of the launch of the Extreme
Ultraviolet Explorer by the Delta II launch vehicle from Cape
Canaveral, Florida. 11,000 orbits later, both the EUVE payload and
the Explorer Platform spacecraft bus continue to perform remarkably
well. The observatory has never been in hardware safehold mode, and
has experienced only one outright hardware failure, the transmit side
of transponder B. EUVE just concluded two important observations, a
pre-comet-impact look at Jupiter, and an observation of the surprise
Target of Opportunity VW-HYI, an outbursting cataclysmic variable star.
SPACECRAFT STATUS
The MMS-based EUVE Explorer Platform continues to perform up to every
expectation. The engineering time planned at the end of last week for
star tracker trend data collection and recalibration of the gyros was
postponed due to the observation of VW-HYI- the unexpected Target of
Opportunity. These activities will be rescheduled at the earliest
practical opportunity; no serious performance degradation will result
from the delay. Another engineering test is being coordinated for
later in the month to characterize the performance of the bright
object sensor of star tracker #1 using the Moon as a probe. This
test is necessary to determine EUVE's observing attitude during the
Jupiter/SL9 impact week.
On day 152, in the process of uploading a new slew table for the
observation of the TOO, an enable bit controlling the articulation of
the solar arrays was inadvertently disabled. This resulted in a period
of reduced charging and subsequently greater than normal discharging
by the batteries during the following orbit night.
Quick action on the part of the Flight Operations Team Attitude
Control System Engineer prevented a much more serious situation,
however, as the arrays were reenabled in time to limit the period of
reduced charging to one orbital day. A review of the steps and
procedures which led to the slew table error is underway.
INSTRUMENT STATUS
No new incidents or problems to report.
UPCOMING EVENTS
V1974-Cygni will be viewed this week, followed by observations of
Zeta-Oph and Comet Mueller.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to operate with only one battery
with two failed cells, while acquiring science data from both
the SAGE and ERBE-NS instruments. All the spacecraft operations
are stable, and the battery is still performing well. ERBS will
perform a frequency test with TDRS this week for the XTE Project.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. A "warm
restart" occurred last week. About 12 "restarts" have occurred
since SAMPEX was launched almost 2 years ago. The last "restart"
was over due, and Code 512 has confirmed that this "restart" has
the same data signature as all the others. These restarts have
little impact to the spacecraft operations or to the scientific
data. However, the ACS now reverts to the old pointing algorithm
for a few hours until the "restart" is detected at the next pass
and until the new ACS software patch is uplinked to the spacecraft.
SAMPEX is currently testing with a new and much smaller JPL
ground station (LEO-D). Twenty-four hours of simultaneous FAST
Campaign and routine SAMPEX operations will be performed in the
SMEX Control Center this week in preparation for the FAST which
is only 2-months from launch.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the PSR 1951+32 science target on June 7.
The next few attitude maneuvers are planned for June 10 to GAL
018+00, June 14 to PSR 1951+32, June 18 to GAL 018+00, and July 5
to PSR 1951+32. These maneuvers are more frequent than usual
because the observations of these targets were broken up to avoid
a star tracker occultation constraint.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of full-time
coverage achieved with 32 Kbps real-time data from May 25 through May
30 was 81 percent, including 18 percent from the GRO Remote Terminal
System (GRTS). An additional 3 percent coverage was achieved with the
BATSE 1 Kbps subset of data during this time. The high-gain antenna
boom angle with respect to the equatorial plane was 7 degrees.
The phase-in period of the GRO Flight Operations Team from TRW to the
NMOS/AlliedSignal contract is proceeding normally toward a July 7,
1994 transition. Of the 21 incumbent technical personnel, 15 have
accepted positions with AlliedSignal.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis.
During the week of 05/30/94 through 06/05/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 05/30/94 through 06/05/94, 7 real
time passes were supported with 26 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
On June 2 the SIGMA 5 was powered down for a building
wide test of the diesel power system. The SIGMA 9 was
on a circuit that was not to lose power during the
test. Just before the test began, the SIGMA 9 crashed
and would not reboot. Because the SIGMA 5 was not
available, the SIGMA 9 had to be reloaded during
operations causing a 44 minute interruption in science
operations. Later in the day, after the power test,
problems were experienced with both computers. The
SIGMA 9 would not lock on telemetry because of problems
with its Telemetry Handling System (THS), and the SIGMA
5 would not send certain commands because it could not
access its disk pack (making science operations
impossible). Only the SIGMA 5 problems may possibly be
linked to the power outage. Because the SIGMA 5
provided good telemetry, and the capability to send
single commands if an emergency arose, it was used for
monitoring purposes until the SIGMA 9 THS was fixed. A
total of 113 minutes of science time was lost in this
instance. In all, 157 minutes of science time were
lost due to ground equipment problems in the OCC.
OPERATIONS
A total of 113 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . 36
Long Wavelength Redundant (LWR) . . . . . . .0
Short Wavelength Primary (SWP) . . . . . . . 77
OPERATIONAL EFFECTIVENESS
Total Time Available : 6720 minutes
Minus losses due to:
Engineering Tests 0 minutes
Spacecraft Problems 0 minutes
Ground System Problems 157 minutes
Operations Problems 0 minutes
Natural Phenomena 0 minutes
Time used in Science Data Ops: 6563 minutes = 97.7%
[email protected]
|
719.15 | May 26 to June 1 | MTWAIN::KLAES | Keep Looking Up | Wed Jun 08 1994 16:37 | 147 |
| From: US4RMC::"[email protected]" "Steve Andrews" 8-JUN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 5/26/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :May 26 to June 1, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
The Extreme Ultraviolet Explorer is operating well; no problems or
unusual incidents to report. Recalibration of the gyros and the
collection of quarterly star tracker trend data (in Survey Mode) are
both planned for later in the week.
SPACECRAFT STATUS
All Explorer Platform systems continue to perform very well. A change
to the charging efficiency of the batteries was made last week in
anticipation of a seasonal dip in the length of spacecraft night. The
strategy of making minor changes in the charging efficiency based on
the length of orbit night is being tested as a new tool in our battery
management efforts. We will carefully study the effectiveness of this
new methodology. Plans for the observation of the Jupiter/SL9
collision continue to be refined. There's a good possibility that moon
presence at the edge of star tracker #1's bright object sensor (BOS)
will not cause the BOS door to close, allowing EUVE's observation of
Jupiter to proceed during periods of the critical impact week when our
planning software indicates the BOS should be closed. This preliminary
conclusion is based upon analysis of previous on-orbit performance
data from the tracker. A special test during June is likely to
confirm this finding and make final plans for July. Plans are also now
finalized between the Flight Operations Team and the GSFC Flight
Dynamics Group to conduct a series of six special slews on June 3rd to
recalibrate EUVE's gyros and collect quarterly star tracker trend data
in Survey Mode.
INSTRUMENT STATUS
No new incidents or problems to report.
UPCOMING EVENTS
A week long observation of EX_HYA will be complete on June 2nd,
followed by twenty eight hours on June 3-4 for gyro recalibration and
star tracker trend data collection.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis was
oriented toward the GAL 25342 science target on May 31. The next
attitude maneuver is planned for June 7 to the PSR 1951+32 target.
Based on the number of packets of data delivered to the instrument
teams from the Packet Processor facility, the percentage of full-time
coverage achieved with 32 Kbps real-time data from May 18 through May
23 was 83 percent, including 19 percent from the GRO Remote Terminal
System (GRTS). An additional 4 percent coverage was achieved with the
BATSE 1 Kbps subset of data during this time. The high-gain antenna
boom angle with respect to the equatorial plane was 16 degrees.
Beginning on May 27, limited amounts of GRO support are being
scheduled routinely on the Second TDRS Ground Terminal to assist in
engineering evaluation.
The phase-in period of the GRO Flight Operations Team from TRW to the
NMOS/AlliedSignal contract is proceeding normally toward a July 7,
1994 transition. Of the 21 incumbent technical personnel, 15 have
accepted positions with AlliedSignal and one more is pending.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis. During
the week of 05/23/94 through 05/29/94, 2 real time passes were
supported with no commands transmitted.
A Command and Telemetry data flow with DTF-93 at JPL was completed
successfully on May 25. The test was to test present and future GCF
configurations.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine basis. During
the week of 05/23/94 through 05/29/94, 7 real time passes were
supported with 35 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
OPERATIONS
A total of 118 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 53
Long Wavelength Redundant (LWR) . . . . . . 4
Short Wavelength Primary (SWP) . . . . . . 61
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to operate with only one battery with
two failed cells, while acquiring science data from both the SAGE and
ERBE-NS instruments. The SAGE instrument has resumed acquiring
science data after being OFF during a 12-day period of continuous
sunlight. The spacecraft battery is still being evaluated after this
long period of continuous sunlight. Extra precautions were taken not
to over-charge the battery in this period because cell failures have
occurred in the past soon after similar periods of continuous sunlight.
The battery charge is now increasing to compensate for the longer eclipse
durations.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data without
any significant problems or operational errors. The ACS change to the
onboard software was uplinked to the spacecraft last week. The Flight
Ops Team (519.1), the Flight Dynamics engineers (553), the GSFC ACS
engineers (712), and the Principal Investigator at the University of
Maryland are all evaluating the results of this new ACS pointing
algorthrum which was requested by the SAMPEX Project Scientist.
Preliminary analysis shows the onboard ACS system is working harder,
which has resulted in slightly larger ACS errors and slightly larger
current spikes onboard the spacecraft. These impacts were all anticipated.
[email protected]
|
719.16 | June 9-15 | MTWAIN::KLAES | Keep Looking Up | Wed Jun 15 1994 15:18 | 129 |
| From: US4RMC::"[email protected]" "Steve Andrews" 15-JUN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, June 9, 1994
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :June 9 to June 15, 1994
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis. During the
week of 06/06/94 through 06/12/94, 3 real time passes were supported with
no commands transmitted.
An ICE telemetry data flow with the Goldstone station D14 to exercise the
new Telemetry Group Controller (TGC-5) was completed successfully on June 9.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine basis. During the
week of 06/06/94 through 06/12/94, 29 real time passes were supported
with commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
OPERATIONS
A total of 140 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 73
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 67
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
Operations of the Explorer Platform and the EUVE science payload were
nominal this week. All flight systems reported good telemetry status.
Observations of the nearby hydrogen filament LVC-88 and Guest Observer
target J1229 are planned for the coming week.
SPACECRAFT STATUS
The Explorer Platform continued its good performance this week. Due to
the shrinking length of orbital day, the battery charging efficiency
value will be changed tomorrow. This will compensate for the slight
undercharging of EUVE's batteries which would otherwise occur. Two TDRS
passes were cut short late last week, resulting in a total of about 24
minutes of real time (recoverable) data loss. The problem passes were
the result of the planning and scheduling effort involved in rejoining
the regular science plan after observing VW-HDI the unexpected Target of
Opportunity of June 2nd. A review of the procedural coordination between
the Flight Operations Team and the Flight Dynamics Group is underway to
prevent future miscues of this kind. Testing of UCB-supplied Absolute
Time Commands (ATCs) for the EUVE payload continued last week by the FOT
on the SDVF simulator. Work is proceeding toward the goal of on-orbit
testing by mid to late July.
INSTRUMENT STATUS
No new incidents or problems to report.
UPCOMING EVENTS
Targets for the coming include LVC-88 and J1229. A special star tracker
test will be performed on the 23rd to characterize tracker #1's ability
to acquire stars while pointing in the vicinity of the bright moon.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis was
oriented toward the GAL 018+00 science target on June 10 and toward PSR
1951+32 on June 14. The next attitude maneuvers are planned for June 18
to GAL 018+00, and July 5 to PSR 1951+32. These maneuvers are more
frequent than usual because the observations of these targets were broken
up to avoid a star tracker occultation constraint.
Based on the number of packets of data delivered to the instrument teams
from the Packet Processor facility, the percentage of full-time coverage
achieved with 32 Kbps real-time data from June 1 through June 6 was 92
percent, including 8 percent from the GRO Remote Terminal System (GRTS).
There was no additional coverage with the BATSE 1 Kbps subset of data
during this time. The high-gain antenna boom angle with respect to the
equatorial plane was 80 degrees.
The phase-in period of the GRO Flight Operations Team from TRW to the
NMOS/AlliedSignal contract is proceeding normally toward a July 7, 1994
transition.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to operate with only one battery with two
failed cells, while acquiring science data from both the SAGE and ERBE-NS
instruments. All the spacecraft operations are stable, and the battery
is still performing well. The space-craft was again successfully yawed
180 degrees this week in order to keep sunlight on the solar arrays.
This was be the 98th time that this maneuver has been performed since the
spacecraft was launched 9.5 years ago.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data without any
significant problems or operational errors. Tests with a new and much
smaller JPL ground station (LEO-D) continue this week, and 900 KBS tests
with the JPL/DSN should start next week. SAMPEX also participated in
multi-mission simulations with the FAST spacecraft last week. The HLT
instrument will continue to be powered OFF/ON every 30-60 days in order
to conserve the isobutane gas within the instrument for the 3-year
extension to the SAMPEX mission.
[email protected]
|
719.17 | June 23-29 | MTWAIN::KLAES | Keep Looking Up | Wed Jun 29 1994 15:22 | 140 |
| From: US4RMC::"[email protected]" "Steve Andrews" 29-JUN-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 6/23/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK :June 23 - June 29, 1994
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis
has been oriented toward the GAL 018+00 science target since June
18. The next attitude maneuver is planned for July 5 to PSR 1951+32.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from June 15 through June 17 was 80 percent, including 18
percent from the GRO Remote Terminal System (GRTS). An
additional 3 percent coverage was achieved with the BATSE 1 Kbps
subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 7 degrees.
The phase-in period of the GRO Flight Operations Team from TRW to
the NMOS/AlliedSignal contract is proceeding normally toward a
July 7, 1994 transition.
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
The Explorer Platform (EP) spacecraft experienced large servo
errors during two consecutive playbacks of Tape Recorders (TR-B)
on Sunday June 26th. Preliminary analysis indicated a short gap
occurred during two successive record cycles where no data was
recorded on the tape. Operational procedures have been modified
to avoid this "bad" section of tape until further analysis can be
performed. A meeting has been scheduled for June 29th to discuss
this unexpected performance.
SPACECRAFT STATUS
On Sunday June 26th a large servo error "spike" was observed
during the 1 Mbps playback of Tape Recorder B (TR-B) during orbit
11381. This data drop-out was reported by PACOR during their
data processing and confirmed by the FOT via realtime plots. The
tape recorder was re-positioned so that the missing data would
not be recorder over and could be captured during the next
playback of TR-B. During the next playback at 512 Kbps the FOT
again observed a large servo error "spike" when the recorder
passed position 80. Processing both the 1 Mbps and 512 Kbps TR-B
dumps revealed an approximate 1 minute and 30 second gap between
tape positions 80 and 83. The same anomaly was observed during
the next playback of TR-B during orbit 11385 which included new
spacecraft data. The FOT has implemented a new tape recorder
management scheme to avoid this "bad" section of tape until
further analysis can be performed.
A Target of Opportunity (TOO) observation was successfully
implemented on Wednesday June 22nd when the dwarf novae SS Cygni,
an eruptive variable star, went into outburst. This unexpected
modification to the weekly observing plan forced cancellation of
a planned Deep Space Network (DSN) test of Transponder B (XPNDR-
B) planned for Friday June 24th. The XPNDR-B test will be
rescheduled at a later date. The TOO was performed around a
special spacecraft engineering test that was conducted on June
23rd to characterize the performance of Fixed Head Star Tracker
(FHST) #1 in the vicinity of the bright moon. This is a
preparatory test of the upcoming observation of the comet
Shoemaker-Levy collision with Jupiter in mid July. Initial
indications are that the tracker will be able to identify it
guide stars even with the nearly full moon being within 12.68
deg. of its boresight.
INSTRUMENT STATUS
The EUVE payload completed the first part of a two part Target of
Opportunity (TOO) observation of the cataclysmic variable star SS
Cyg. The TOO will be re-visited again later this week after
completion of the Guest Observer target LVC-88-+36. The EUVE
instrument continues to work well and there are no new incidents
or problems to report.
UPCOMING EVENTS
Perform background observation of Jupiter prior to the impact of comet
Shoemaker-Levi. Accommodate the June 30th leap second as necessary.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis.
During the week of 06/21/94 through 06/27/94, 1 real time
pass was supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 06/21/94 through 06/27/94, 7 real
time passes were supported with 41 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
Two instances of data and command problems associated
with the recent switch to the FTS 2000 line occurred
this past week. A total of 7 minutes of science time
was lost due to these problems.
OPERATIONS
A total of 144 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 39
Long Wavelength Redundant (LWR) . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . .105
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . 35 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6713 minutes
Effectiveness. . . . . . . . . . . . . . . .99.9%
[email protected]
|
719.18 | June 30 to July 6 | MTWAIN::KLAES | Houston, Tranquility Base here... | Wed Jul 06 1994 16:16 | 113 |
| From: US2RMC::"[email protected]" "Steve Andrews" 6-JUL-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 6/30/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK :June 30 to July 6, 1994
COMPTON GAMMA RAY OBSERVATORY R. Wilson
The Compton observatory is in normal pointing mode. The +Z
axis was oriented toward the PSR 1951+32 science target on July 5. The
next attitude maneuver is planned for July 12 to GAL 253-42.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time data
from June 19 through June 27 was 84 percent, including 18 percent from
the GRO Remote Terminal System (GRTS). An additional 3 percent coverage
was achieved with the BATSE 1 Kbps subset of data during this time. The
high-gain antenna boom angle with respect to the equatorial plane was 47
degrees.
The phase-in period of the GRO Flight Operations Team from TRW to the
NMOS/AlliedSignal contract is proceeding normally toward a July 7, 1994
transition.
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
EUVE is currently conducting an important pre-comet-impact
observation of the planet Jupiter. All payload and spacecraft
systems are operating well. A gyro recalibration and observations of
M15/AC211 and the star RE2009 are scheduled for later this week.
SPACECRAFT STATUS
Performance of the Explorer Platform has been nominal this week.
Procedural changes have been implemented to avoid the section of tape on
recorder B which caused data errors last week. Research on the nature of
the problem is still underway.
Observation of the planet Jupiter this week has required special power
management steps to be taken due to its high antisun location and the
required alignment (spacecraft roll) of the spectrometer with respect to
the Jupiter/Io system. The spacecraft is currently flying with a 20
degree solar array offset with respect to the sun. Additionally, no
transition to constant current mode (CCM) at beginning of orbit day is
being commanded. Transition to our normal 40 degree offset, with CCM at
orbit enabled at orbit dawn will be effected with the observation of M15.
A gyro recalibration consisting of six preplanned slews, and a brief
period of Survey Mode to gather star tracker trend data are planned for
July 8, 9. The first on orbit test of UCB-supplied absolute time commands
of the EUVE payload should also occur later this week. A series of tests
to prove proper operation of the stored commands as well as their
response to inhibit conditions will be performed.
INSTRUMENT STATUS
No new problems or incidents to report.
UPCOMING EVENTS
Gyro recalibration and a brief period in Survey Mode for star
tracker trend data collection are planned for July 8-9.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis. During the
week of 06/26/94 through 07/03/94, 2 real time passes were supported with
no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 06/26/94 through 07/03/94, 7
real time passes were supported with 25 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
On June 30, forty minutes of science time were lost
when the Sigma 5 computer experienced problems
accessing its shared disk pack. A switch was made to
the Sigma 9 computer and operations resumed.
OPERATIONS
A total of 139 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 57
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 82
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . 40 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6680 minutes
Effectiveness. . . . . . . . . . . . . . . .99.4%
[email protected]
|
719.19 | July 7-13 | MTWAIN::KLAES | Houston, Tranquility Base here... | Wed Jul 20 1994 13:45 | 195 |
| From: US4RMC::"[email protected]" "Steve Andrews" 13-JUL-1994
To: [email protected]
CC:
Subj: Orbiting satellite, 7/7/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :July 7 to July 13, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
EUVE conducted a week of nominal operations which concluded with
a short engineering test period. Currently we are taking a
dithered observation of the star RE2009. All systems are green.
Observation of the planet Jupiter in anticipation of its impact
with comet Shoemaker/Levy will begin on July 13th.
SPACECRAFT STATUS
The Explorer Platform Spacecraft bus continues to perform well.
The first on-orbit executions of payload commands from the
spacecraft's Absolute Time Processor (ATC) were successfully
performed last week. Incrementally more complex ATC tasks will
be performed this week and in the weeks to come until this mode
of operation becomes standard for the mission. Tests of the
process to inhibit the execution of the payload ATCs- desirable
during instrument anomalies- were also successful.
Six specially designed slews were conducted over the weekend as
part of the process of gyro recalibration being conducted through
the analysis and support of the GSFC Flight Dynamics Facility
(FDF). The spacecraft also spent twelve hours in Survey Mode
to gather quarterly performance trend data on EUVE's star trackers.
EUVE will conduct a scientifically important observation of the
collision of Comet Shoemaker/Levy-9 with the planet Jupiter this
week. Much engineering planning and testing has been
accomplished to define the operations necessary to achieve the
maximized science for this event. EUVE will roll to an attitude
outside its normal constraints to begin the observation, and will
slowly increase its antisun pitch until this too violates the
defined guidelines. Because of the high pitch and roll, careful
management of spacecraft power will be necessary.
INSTRUMENT STATUS
No new problems or incidents to report.
UPCOMING EVENTS
EUVE will observe the effects of the comet collision on Jupiter's
Magnetosphere and IO torus throughout much of the next few weeks.
A short observation of the star GD-246 will take place on July
16/17 when moon presence in the Deep Survey Spectrometer prevents
pointing towards Jupiter.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GAL 253-42 science target on July 12.
The next attitude maneuver is planned for July 18 to GAL 009-08.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from June 28 through July 4 was 82 percent, including 17
percent from the GRO Remote Terminal System (GRTS). An
additional 3 percent coverage was achieved with the BATSE 1 Kbps
subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 47 degrees.
On July 8, at the request of the HST Project, the GRO transmitter
was turned off during two TDRS contacts to prevent the GRO signal
from interfering with the HST while HST was recovering from a
safe mode. Interference was being experienced by HST because
the HST low-gain antenna was in use while HST was in a safe mode.
After the HST recovery, there was no further interference.
Approximately 1.8 hours of GRO data were lost while the
transmitter was off.
The transition of the GRO Flight Operations Team from TRW to the
NMOS/AlliedSignal contract was made as planned on July 7.
Fifteen of the 21 TRW engineers and analysts converted to
AlliedSignal to stay with the GRO Project. Six experienced
AlliedSignal engineers and analysts were trained in GRO
operations to round out the team. The three-month phase-in
period was completed very smoothly. The Project would like to
thank all those involved in this very successful implementation.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to acquire science data from both
the SAGE and ERBE-NS instruments while operating with only one
battery which has two failed cells. All the spacecraft operations
remain stable, and the battery is still performing well. A number
of STGT tests at all data rates were successfully completed last
week. A spacecraft yaw maneuver will be performed tomorrow, using
a slightly modified procedure.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. Tests
with a new and much smaller JPL ground station (LEO-D) are
continuing, as well as 900 KBS "strip & ship" tests with the
JPL/DSN stations. Tests are also being performed with the Poker
Flat station in Alaska as part of the preparation for the FAST
Launch. The experimenter reported that one of the detectors in
the MASS instrument has become quite noisy, and the entire
instrument will be turned OFF for four days to see if it
recovers. The FAST and SAMPEX Flight Ops Teams have now combined
and they are all receiving intensive training in preparation for
the FAST launch.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Support of the ICE spacecraft continues on a routine basis.
During the week of 07/04/94 through 07/10/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 07/04/94 through 07/10/94, 7 real
time passes were supported with 31 commands transmitted.
A IMP-8 command test was conducted on July 8 to check out a
new prototype Network Command Processing System (NCPS) with
the Wallops Island Station. At the direction of the IMP
project only standard routine commands were transmitted.
The test commands were verified, and the station will be
using the NCPS vice the Spacecraft Command Encoder (SCE) to
command the spacecraft for an indefinite period.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
HIGHLIGHTS
On July 6, the VILSPA station experienced problems with
their ground system which prevented them from taking control
of the spacecraft at the specified time. In order to avoid
a waste of potential science time, IUESOC and IUEOCC
performed the operations requested by the VILSPA operators
until they were ready to assume control of the spacecraft.
On July 7 patches to the onboard computer code needed for a
special control mode used during earth eclipse periods were
uplinked to the spacecraft. IUE's 34th shadow season begins
on July 14 and runs through August 8.
An unexplained 11 minute loss of data occurred during a
camera read on July 8. Although this data loss lasted for
11 minutes, a total of 72 minutes of science time was
wasted. This 72 minutes was the time required to prepare
and expose the camera.
A spacecraft anomaly, V-60, occurred on July 10. In this
anomaly, communications between the spacecraft decoders and
the onboard computer are interrupted, resulting in a loss of
command data being transferred into the onboard computer's
memory. This anomaly has occurred several times in the past.
OPERATIONS
A total of 124 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 48
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 76
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . 72 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6648 minutes
Effectiveness. . . . . . . . . . . . . . . .98.9%
[email protected]
|
719.20 | July 14-20 | MTWAIN::KLAES | Houston, Tranquility Base here... | Thu Jul 21 1994 14:09 | 146 |
| From: US4RMC::"[email protected]" "Steve Andrews" 21-JUL-1994
To: [email protected]
CC:
Subj: Orbiting Satellites 7/14/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :July 14 to July 20, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
Operations with the Extreme Ultraviolet Explorer went smoothly during the
reporting period; both the spacecraft and payload are functioning very
well. Observations of the planet Jupiter began midweek and will continue
with a few short interruptions until August 12.
SPACECRAFT STATUS
To maintain consistent battery charging levels and profiles, constant
current mode at orbital daybreak was disabled and the solar arrays moved
to a 20 degree solar offset angle during the past week. Both these steps
were mandated by the high antisun and roll angles of the spacecraft
needed to observe the Jupiter/comet impact. A change in battery charging
efficiency was also commanded to account for the increasing length of
EUVE's orbital day. Careful battery management has resulted in
consistently good performance from the three batteries. All other
subsystems are operating nominally. The current Shuttle mission has not
had a negative impact on the SCE or HGA gimbal temperatures, despite
EUVE's use of longer TDRS tracks.
Testing of UCB-supplied Absolute Time Commands (ATC) continued during the
past week. A software bug identified in one of the GSFC ground
processing systems should be fixed in the coming week which will allow
testing to continue. Regular telecons with Berkeley have been held on
this subject.
INSTRUMENT STATUS
No new problems or incidents to report.
UPCOMING EVENTS
EUVE will observe the effects of the comet collision on Jupiter's
Magnetosphere and Io torus throughout much of the next few weeks.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
The Compton observatory is in normal pointing mode. The +Z axis was
oriented toward the GAL 009-08 science target on July 18. The next
attitude maneuver is planned for July 25 to GAL 253-42. Based on the
number of packets of data delivered to the instrument teams from the
Packet Processor facility, the percentage of full-time coverage achieved
with 32 Kbps real-time data from July 6 through July 11 was 78 percent,
including 18 percent from the GRO Remote Terminal System (GRTS). An
additional 4 percent coverage was achieved with the BATSE 1 Kbps subset
of data during this time. The high-gain antenna boom angle with respect
to the equatorial plane was 7 degrees.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
Preparations for the July 19th ICE/IMP move are being made.
Two ICE realtime passes were supported this week with no commands
transmitted to the spacecraft.
The ICE/IMP relocation to IUE should occur on 19 July.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
Six IMP realtime passes were supported this week with twenty commands
transmitted to the spacecraft. The data lines from Nascom through the
TOR to the Nascom line drivers in room W10D (the new TPOCC location)
have been checked during realtime support and indicate data is being
received.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
A pre-shadow season battery discharge was performed on July 13. The
purpose of this procedure was to check out the health of the batteries
and to also begin the rejuvenation process which occurs when the
batteries are discharged and recharged multiple times.
IUE's 34th shadow season began on July 14. To date the spacecraft has
performed nominally.
OPERATIONS
A total of 173 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . 69
Long Wavelength Redundant (LWR) . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . 104
OPERATIONAL EFFECTIVENESS
Total Time Available : 6720 minutes
Minus losses due to:
Engineering Tests...............................0 minutes
Spacecraft Problems.............................0 minutes
Ground System Problems..........................0 minutes
Operations Problems.............................0 minutes
Natural Phenomena...............................0 minutes
Time used in Science Data Ops:...............6720 minutes
Effectiveness.......................................100 %
EQUIPMENT STATUS
All equipment is currently operational.
PROBLEMS AND CONCERNS
None
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
The ERBS spacecraft continues to acquire science data from both the
SAGE and ERBE-NS instruments while operating with only one battery
which has two failed cells. All the spacecraft operations remain
stable, and the battery is still performing well. A spacecraft yaw
maneuver was successfully performed last week. There were no special
operations or network tests this past week.
SAMPEX - J. Williamson
SAMPEX continues to acquire all science and engineering data without any
significant problems or operational errors. Tests with a new and much
smaller JPL ground station (LEO-D) are continuing, as well as tests with
the Poker Flat station in Alaska. The Leica instrument turned itself OFF
several times this week when its HV noise became too high. The MASS
instrument was also turned OFF for four days this week in order to reduce
the noise in one of its detectors. The HLT instrument was turned back ON
after being OFF for about one month. The HLT experimenter is turning OFF
his instrument periodically in order to conserve the isobutane gas within
the instrument. The HLT instrument will stay ON for about two months.
[email protected]
|
719.21 | July 21-27 | MTWAIN::KLAES | Houston, Tranquility Base here... | Tue Jul 26 1994 18:04 | 168 |
| From: US4RMC::"[email protected]" "Steve Andrews" 26-JUL-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 7/21/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK :July 21 to July 27, 1994
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------
On July 19 problems with the Sigma 5 Telemetry Handling
System (THS) resulted in a loss of 17 minutes of science
time. The Sigma 5 THS failed during operations, requiring a
switch to the Sigma 9 before operations could resume.
IUE's 34th shadow season is in progress. To date the
batteries have performed as expected and all operations
associated with each of the daily eclipse periods have been
performed without any problems.
OPERATIONS
A total of 162 images were taken during the week:
Long Wavelength Primary (LWP) ................. 65
Long Wavelength Redundant (LWR)..................0
Short Wavelength Primary (SWP) .................97
OPERATIONAL EFFECTIVENESS
Total Time Available : 6720 minutes
Minus losses due to:
Engineering Tests 0 minutes
Spacecraft Problems 0 minutes
Ground System Problems 17 minutes
Operations Problems 0 minutes
Natural Phenomena 0 minutes
Time used in Science Data Ops: 6703 minutes = 99.7 %
EQUIPMENT STATUS
All equipment is currently operational.
PROBLEMS AND CONCERNS
None.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
--------------------------------
The ICE/IMP move to IUE was completed on 20 July.
Packet Processor System Release 3.1 was delivered on 18
July. Acceptance testing is in progress.
One ICE realtime passes were supported this week with
no commands transmitted to the spacecraft.
The capability of the Flight Operations Team to command
the ICE spacecraft after the move to IUE is still being
investigated.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
-----------------------------------
Seven IMP realtime passes were supported this week with
twenty-two commands transmitted to the spacecraft.
EXTREME ULTRAVIOLET EXPLORER (EUVE) -K. Hartnett
-----------------------------
EUVE continues to perform admirably; no new payload or
spacecraft problems to report. Jupiter was EUVE's target
for virtually all of last week, as comet Shoemaker/Levy 9
slammed into the Jovian atmosphere. Early results from
EUVE include the post-impact detection of the element
Helium. No changes have yet been seen in the Io torus.
The star Alpha Erindani will be observed for three days
this week before a return look at Jupiter on the 29th.
SPACECRAFT STATUS
All spacecraft subsystems performed nominally this week.
A nonstandard power arrangement due to the high spacecraft
roll was implemented prior to the observation of Jupiter.
Battery response in general has been good. Battery #3 is
showing some signs of stress and is currently at a
differential voltage value of 100 mV. This level will
most likely decline with the return of our standard power
configuration later in the week.
One realtime pass was lost this week (data recoverable)
due to the inadvertent deletion of a high gain antenna
track by a member of the Flight Operations Team. New
console procedures have been implemented to guard against
such occurrences in the future. Testing of UCB-supplied
ATC commands continues with steady progress. This week's
testing will center around specific command macros that
UCB intends to use on orbit.
An opportunity to perform a transponder B residual output
test with the DSN on DOY 202/203 was not realized because
JPL could not support the testing. New opportunities will
be sought to fulfill this request of the GSFC Transponder
Failure Review Team.
INSTRUMENT STATUS
No new problems or incidents to report. Scanner C was
turned off, the short wave spectrometer's count rate
shutdown value was increased, and the unused sections of
the Deep Survey detector were blanked to maximize the
science data bandwidth for the Jupiter/comet observation.
UPCOMING EVENTS
EUVE will observe Alpha Erindani, then Jupiter, then the
star Wolf 630 in the week to come.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------
The Compton observatory is in normal pointing mode. The +Z
axis was oriented toward the GAL 253-42 science target on
July 25. The next attitude maneuver is planned for August 1
to GAL 088-47.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps
real-time data from July 13 through July 17 was 92 percent,
including 9 percent from the GRO Remote Terminal System
(GRTS). No additional coverage was achieved with the BATSE
1 Kbps subset of data during this time. The high-gain
antenna boom angle with respect to the equatorial plane was
70 degrees.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
---------------------------------
The ERBS spacecraft continues to acquire science data from
both the SAGE and ERBE-NS instruments while operating with
only one battery. All the spacecraft operations remain
stable, and the battery continues to perform well.
SAMPEX - J. Williamson
----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors.
Tests with a new and much smaller JPL ground station (LEO-D)
are continuing, as well as tests with the DSN stations. The
SAMPEX Flight Ops Team (FOT) is still cross training with
the FAST FOT in preparation for the FAST launch.
[email protected]
|
719.22 | Conference on Space Applications of AI - May 1995 | MTWAIN::KLAES | No Guts, No Galaxy | Wed Aug 03 1994 13:23 | 89 |
| From: US4RMC::"[email protected]" "James Rash" 3-AUG-1994 02:03:52.88
To: [email protected]
CC:
Subj: CFP: 1995 Goddard Conference on Space Applications of AI/IT.
Call for Papers
1995 Goddard Conference on
Space Applications of Artificial Intelligence and
Emerging Information Technologies
May, 1995
NASA Goddard Space Flight Center
Greenbelt, Maryland
The Tenth Annual Goddard Conference on Space Applications of
Artificial Intelligence and Emerging Information Technologies
will focus on AI and IT research and applications relevant to
space systems, space operations, and space science.
Topics will include, but are not limited to:
o Neural networks, genetic algorithms & fuzzy systems
o On-orbit process control, monitoring, and analysis
o Knowledge-based spacecraft command & control
o High-Performance Computing & Communications
o Expert system management & methodologies
o Distributed knowledge-based systems
o Fault-tolerant/autonomous systems
o Intelligent database management
o Robotics & telerobotics
o Planning & scheduling
o Knowledge acquisition
o Virtual reality
o Image analysis
Original, unpublished papers are now being solicited for the
conference. Abstracts should be 300-500 words in length, and
must describe work with clear AI or IT content and
applicability to space-related problems. Two copies of the
abstract should be submitted by September 2, 1994, along with
the authors name, affiliation, address, email address (if
available), and telephone number. Notification of tentative
acceptance will be given by September 19, 1994. Papers
should be no longer than 15 pages and must be submitted in
camera-ready form for final acceptance by November 25, 1994.
Accepted papers will be presented formally or as poster
presentations, which may include demonstrations. All
accepted papers will be published in the conference
proceedings as an official NASA document, which will be
distributed at the time of the conference. There will be a
conference award for Best Paper.
No commercial presentations will be accepted.
Send abstracts to:
Walt Truszkowski
NASA/GSFC
Code 522.3
Greenbelt, MD 20771
[email protected]
FAX: 301-286-1768
Please FAX abstracts, or send via email, if possible.
Important dates:
Abstracts due: Sept. 2, 1994
Papers due Nov. 25, 1994
The conference call for papers, announcements, and other
information can be accessed via WORLD WIDE WEB at
http://ddwilson.gsfc.nasa.gov.
For further information, call (301) 286-3150.
This conference is sponsored by the Mission Operations and Data
Systems Directorate, Code 500, NASA Goddard Space Flight Center.
Please forward and/or print and post copies of this call for papers.
James Rash
NASA/GSFC, Code 531.1
Greenbelt, Maryland 20771
USA
Phone: 301-286-3595 (voice)
301-286-1724 (fax)
|
719.23 | July 28 to August 3 | MTWAIN::KLAES | No Guts, No Galaxy | Wed Aug 03 1994 18:05 | 164 |
| From: US4RMC::"[email protected]" "Steve Andrews" 3-AUG-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, July 28 to Aug. 3, 1994
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK :July 28 to Aug. 3, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
-----------------------------------
Operations with the Extreme Ultraviolet Explorer went smoothly
this week. There were no unexpected payload or spacecraft
performance anomalies. A return look at the Jovian system was
completed over the weekend. Next week's observing time is
devoted to the M dwarf star, Wolf 630.
SPACECRAFT STATUS
All spacecraft subsystems performed nominally this week. The
power subsystem was reconfigured several times as a result of the
science plan taking us from Jupiter to the star Alpha Erindani
and then back to Jupiter again. The Jupiter observations
required special configurations due to their high antisun and
spacecraft roll requirements. All three batteries have shown
small differential voltage changes as a result of the
reconfigurations, but these were expected and, given their small
magnitude, are of no present concern. ATC load testing also
continues. The most recent load had four macros with a total of
563 commands. A new GSFC CMS delivery and an EPPS fix should
improve the overall testing process in the coming week.
INSTRUMENT STATUS
No new problems or incidents to report. Scanner C was turned
off, the short wave spectrometer's count rate shutdown value was
increased, and the unused sections of the Deep Survey detector
were blanked to maximize the science data bandwidth for the
Jupiter/comet observation.
UPCOMING EVENTS
EUVE will observe the M dwarf star Wolf 630 until August 7th.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------------
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GAL 088-47 science target on August 1.
The next attitude maneuver is planned for August 8 to PKS 0528+134.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from July 19 through July 24 was 86 percent, including 14
percent from the GRO Remote Terminal System (GRTS). An
additional 1 percent coverage was achieved with the BATSE 1 Kbps
subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 59 degrees.
Limited amounts of GRO support are continuing to be scheduled
routinely on the Second TDRS Ground Terminal to assist in
engineering evaluation. Beginning on August 3, the support will
be increased to approximately two contacts per day.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
---------------------------------------
The ERBS spacecraft continues to acquire science data from both
the SAGE and ERBE-NS instruments while operating with only one
battery. However, the SAGE instrument is currently OFF while the
spacecraft is a period of continuous sunlight for approx. 10 days
(July 30 - Aug 10) because there are no SAGE science events
during this time. Battery Ops are also very critical during this
time because of the surplus of power onboard the spacecraft could
quickly overcharge the battery.
SAMPEX - J. Williamson
-----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. Tests
with a new and much smaller JPL ground station (LEO-D) are
continuing for the FAST Project, as well as ranging tests with
the DSN stations for the RADARSAT Project. The combined
FAST/SAMPEX Flight Ops Team (FOT) is now down-sizing as a result
of the 1-year delay in the FAST launch. The Space Power
Applications Branch (734) has requested the battery charge be
reduce to level-4 on SAMPEX. The SAMPEX FOT is now evaluating
this request and joint meetings will be held this week.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 07/25/94 through 07/31/94, 2 real time
passes were supported with 5 commands transmitted.
The Ogilvie experiment (OGH) was turned back On, on July 28,
as the spacecraft power has increased as it travels closer
to the Sun.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 07/25/94 through 07/31/94, 6 real
time passes were supported with 25 commands transmitted.
ICE/IMP TPOCC
--------------
The ICE/IMP TPOCC is now located in room W10D in Building
14. The first ICE pass supported in the new location was at
1125Z on July 27, 1994. The first ICE command pass was on
July 28, 1994. Both passes were successfully completed.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
Three instances of an anomaly in which a serial command
is not properly transferred from the spacecraft command
decoder to the specified device occurred this week.
This type anomaly has occurred throughout the history
of IUE; however, due to a lack of relevant data in
telemetry, it is impossible to pinpoint the cause.
IUE's 34th shadow season is progressing as expected.
The last daily eclipse period will occur on Aug. 7.
OPERATIONS
A total of 131 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 55
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 76
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
[email protected]
|
719.24 | August 4-10 | MTWAIN::KLAES | No Guts, No Galaxy | Tue Aug 16 1994 14:45 | 113 |
| From: US1RMC::"[email protected]" "Steve Andrews" 16-AUG-1994
To: [email protected]
CC:
Subj: Orbiting satellites, 8/4/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :Aug. 4 to Aug. 10, 1994
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 08/01/94 through 08/07/94, 1 real time
pass was supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 08/01/94 through 08/07/94, 7 real
time passes were supported with 26 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------
HIGHLIGHTS
IUE's 34th shadow season completed on Aug 7. This
shadow season progressed as anticipated and the battery
performance is now being analyzed in preparation for
next year's shadow seasons.
OPERATIONS
A total of 129 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 59
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 70
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
COMPTON GAMMA RAY OBSERVATORY R. Wilson
-----------------------------
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GRO J1655-40 on August 4 to observe an
x-ray nova in Scorpius as a target of opportunity. The observation
of the target of opportunity was completed with a maneuver to the
previously scheduled PKS 0528+134 science target on August 9.
The next attitude maneuver is planned for August 29 to CYG X-3.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from July 26 through July 31 was 87 percent, including 12
percent from the GRO Remote Terminal System (GRTS). Additional
coverage of less than 1 percent was achieved with the BATSE 1
Kbps subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 70 degrees.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
------------------------
The ERBS spacecraft continues to acquire science data from the
ERBE-NS instrument. The SAGE instrument is currently OFF while
the spacecraft is a 10-day period of continuous sunlight (July 30
Aug 10). The SAGE instrument requires orbital sunrises and
sunsets for their scientific data. Battery Ops are very critical
during this full-sun period because the surplus of power from the
solar arrays could quickly overcharge the battery.
SAMPEX - J. Williamson
-----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. Tests
with a new and much smaller JPL ground station (LEO-D) continue,
as well as 900 KBS tests with the DSN stations and operational
tests with the Poker Flat station in Alaska. There were no
operational changes in the instruments this past week. However,
the battery charge level was reduced at the request of the GSFC
battery engineers (Code 734). They were trying to reduce some
spikes in the battery differential voltage which occur at the
beginning of each charge cycle. However, the battery charge level
was returned to its original level after a few days when the
voltage spikes did not decrease as expected. The Code 734 power
engineers are studying the test results and will probably try
something else in 1-2 weeks after the eclipse durations have
decreased further.
[email protected]
|
719.25 | August 11-17 | 30254::KLAES | No Guts, No Galaxy | Fri Aug 19 1994 15:38 | 164 |
| From: US1RMC::"[email protected]" "Steve Andrews" 17-AUG-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 8/11/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :Aug. 11 to Aug. 17, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
-----------------------------------
EUVE continues to perform well, with no new hardware or software
incidents to report on either the telescopic payload or
spacecraft bus. EUVE's post-comet-collision observation of the
planet Jupiter was completed this week, with the observatory
pointing at a new mission-high antisun angle of 102 degrees.
Observing time in the coming week is devoted to the hot star EU-2214.
SPACECRAFT STATUS
The EUVE Explorer Platform is performing robustly. Transition
back to the spacecraft's normal power configuration was
accomplished without incident this week following the high
antisun (102 degree) observation of the planet Jupiter. Battery
performance has been nominal. Twenty six serial magnitude
absolute time commands were successfully loaded into and executed
on-orbit by the CDP on 8/12. This activity was part of the
ongoing effort to implement a fully stored-command mode of
operation for the EUVE payload. Ground testing of these loads in
the GSFC SDVF simulator also recently established the maximum
number of commands that can execute on a single time tag. This
information now permits UCB to begin packaging their routine
command activities into additional stored command lists which
will be tested in the weeks to come. A string of negative
acquisitions occurred on August 11 during the second day of the
EUVE-supported Second TDRS Ground Terminal (STGT) 48-hour loading test.
An improper ephemeris product from FDF (with TDRS West and Spare
reversed in the table) were at the root of the problem, which took the
better part of a shift to fully resolve. Careful and judicious tape
recorder management by the Flight Operations Team, however, prevented
any recorded science data loss. A DSN Canberra pass on 8/9 failed to
record any output from EUVE's transmitter "B" which stopped radiating
on April 2nd. Periodic checks of the transmitter will be conducted in
the future.
INSTRUMENT STATUS
No new problems or incidents to report. During the Jupiter
observation Scanner C was turned off, the short Wave spectro-
meter's count rate shutdown value was increased, and the unused
sections of the Deep Survey detector were be blanked to maximize
the available science data bandwidth.
UPCOMING EVENTS
"New Opacity, Model Atmospheres and the Abundance Pattern in Hot
DA Stars" is the title of the proposed science resulting from
EUVE's observation of EU_2214 in the week to come.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 08/08/94 through 08/14/94, 1 real time
pass was supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 08/08/94 through 08/14/94, 2 real
time passes were supported with 29 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
The patches to the onboard computer code required for
shadow season were removed on Aug 9. This operation
did not interfere with science operations and no
problems were encountered.
Ten minutes of science time were lost on Aug 12 due
power hits. Following a couple of power hits the
shared disk pack memory unit being used would not
function properly and had to be switched with another unit.
OPERATIONS
A total of 141 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 59
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) .. . . . . . . 82
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests . . . . . . . . .0 minutes
Spacecraft Problems . . . . . . . .0 minutes
Ground System Problems. . . . . . 10 minutes
Operations Problems . . . . . . . .0 minutes
Natural Phenomena . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6710 minutes
Effectiveness. . . . . . . . . . . . . . . .99.9%
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------------
The Compton observatory is in normal pointing mode. The +Z axis
remains oriented toward the PKS 0528+134 science target. The
next attitude maneuver is planned for August 29 for a two-day
return observation of the GRO J1655-40, an x-ray nova in Scorpius
which was observed as a target of opportunity between August 4-9.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from August 2 through August 3 was 80 percent, including 16
percent from the GRO Remote Terminal System (GRTS). Additional
coverage of 3 percent was achieved with the BATSE 1 Kbps subset
of data during this time. The high-gain antenna boom angle with
respect to the equatorial plane was 28 degrees. For the target
of opportunity, the percentage of coverage achieved with 32 Kbps
real-time data from August 5 through August 8 was 82 percent,
including 20 percent from the GRTS. Additional coverage of 1
percent was achieved with the BATSE 1 Kbps subset of data. The
high-gain antenna boom angle with respect to the equatorial plane
was 15 degrees.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
---------------------------------------
The ERBS spacecraft continues to acquire science data from the
ERBE-NS and the SAGE instruments. The 10-day period of continuous
sunlight ended this past week. The battery, the spacecraft and
the instruments operations are all normal.
SAMPEX - J. Williamson
-----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. Another
software "warm restart" occurred this past week during a
real-time pass. This was the twelfth one since the SAMPEX was
launched over two years ago, and 2.5 months since the last one.
There was no data loss as a result of the "warm restarts",
because the spacecraft automatically recovers. However, the
Flight Ops Team (FOT) must manually reload some contingency
procedures and some software patches everytime one of these "warm
restarts" occur. There were no other changes in either the
spacecraft or instrument operations this past week.
[email protected]
|
719.26 | August 18-24 | MTWAIN::KLAES | No Guts, No Galaxy | Wed Aug 24 1994 11:12 | 147 |
| From: US1RMC::"[email protected]" "Steve Andrews" 24-AUG-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 8/18/94 to 8/24/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :Aug. 18 to Aug. 24, 1994
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 08/15/94 through 08/21/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 08/15/94 through 08/21/94, 7 real
time passes were supported with 25 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
New data and voice lines between NASCOM and the VILSPA
station were tested on August 17. The switch to these
new lines is necessary due to the planned elimination
of the Madrid switching station. The two data lines
used for the ISLU (Interface Switching and Logic Unit)
emergency data and command functions performed
nominally. The primary voice line performed nominally
however, the backup voice line experienced problems.
The problem with the backup voice line is still being
investigated.
OPERATIONS
A total of 92 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 44
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 48
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .10 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6710 minutes
Effectiveness. . . . . . . . . . . . . . . 99.85%
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
-----------------------------------
The Extreme Ultraviolet Explorer was dedicated to the observation
of the hot star EU-2214 during the past week. All spacecraft and
payload subsystems are working nominally. The scheduled targets
for the coming week are HR-1099, the Moon, and CF-TUC.
SPACECRAFT STATUS
The EUVE Explorer Platform is performing well. Battery
performance is being closely monitored since the use of Constant
Current Charge Mode at orbit dawn has been discontinued. So far
performance has been as expected.
The GSFC Flight Operations Teams continues to support the Space
Networks' testing of the Second TDRSS Ground Terminal (STGT). On
Friday, August 26th, a two week Navigation Certification Test
will get underway. Testing will also continue this week on the
use of payload ATC commanding.
INSTRUMENT STATUS
No new problems or incidents reported by UCB.
UPCOMING EVENTS
EUVE will interrupt the observation of HR-1099 late on Wednesday
to collect one orbit's worth of data on the waning gibbous Moon.
On the 28th, EUVE slews to the southern constellation of the
Toucan (Tucana) to observe the star CF-TUC.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------------
The Compton observatory is in normal pointing mode. The +Z axis
remains oriented toward the PKS 0528+134 science target. The
next attitude maneuver is planned for August 29 for a two-day
return observation of the GRO J1655-40, an x-ray nova in Scorpius
which was observed as a target of opportunity between August 4
and August 9. An attitude maneuver toward the Vela Pulsar is
planned for August 31.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from August 10 through August 15 was 83 percent, including
19 percent from the GRO Remote Terminal System (GRTS).
Additional coverage of 3 percent was achieved with the BATSE 1
Kbps subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 23 degrees.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
---------------------------------------
The ERBS spacecraft continues to acquire science data from both
the ERBE-NS and the SAGE instruments. The spacecraft experienced
a "Hit" in its command memory this past week; however, there was
no impact. The 100th spacecraft yaw-turn will be performed next
week when the spacecraft eclipses are of maximum duration.
SAMPEX - J. Williamson
-----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The C/D
ratio of the battery charger was decreased slightly this past
week at the request of the GSFC battery engineers (Code 734).
They are now evaluating the results of this change on the battery
charging profile as the duration of the eclipses continue to
decrease. The spacecraft will enter a period of continuous
sunlight in early September.
[email protected]
|
719.27 | August 25-31 | MTWAIN::KLAES | No Guts, No Galaxy | Wed Sep 07 1994 19:23 | 113 |
| From: US1RMC::"[email protected]" "Steve Andrews" 1-SEP-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 8/25/94 - 8/31/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :Aug. 25, 1994 to Aug. 31, 1994
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 08/22/94 through 08/28/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 08/22/94 through 08/28/94, 7
real time passes were supported with 48 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
On Aug 28 the ground system skipped a line of a
procedure and the expected command was not sent to the
spacecraft. The procedure was re-executed immediately
without incident.
Observations of Venus took place on Aug 28 and Aug 29.
These were power negative observations at a Sun angle
of 134 degrees using Santiago as a backup station. All
operations were nominal and proceeded without incident.
OPERATIONS
A total of 195 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 92
Long Wavelength Redundant (LWR) . . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . . 103
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------------
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the GRO J1655-40 science target on August 29
for a two-day return observation of this x-ray nova in Scorpius
which was discovered by the GRO BATSE instrument and observed as
a GRO target of opportunity between August 4 and August 9. An
attitude maneuver toward the Vela Pulsar is planned for August 31.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from August 16 through August 22 was 82 percent, including
18 percent from the GRO Remote Terminal System (GRTS).
Additional coverage of 5 percent was achieved with the BATSE 1
Kbps subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 23 degrees.
EARTH RADIATIONS BUDGET SATELLITE (ERBS) - J. Williamson
----------------------------------------
The ERBS spacecraft continues to acquire science data from both
the ERBE-NS and the SAGE instruments. Their battery and other
operations remain stable. The one-hundredth spacecraft yaw-turn
was successfully performed last week. The spacecraft also
experienced another "bit flip" in its command memory this past
week; however, there was no impact.
SAMPEX - J. Williamson
-------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
voltage level of the battery charger was decreased last week at
the request of the GSFC battery engineers (Code 734). They are
now evaluating the results of this change on the battery charging
profile, and to determine if more changes are required as the
eclipse durations continue to decrease. Changes were also made to
the onboard Data Processing Unit (DPU) at the request of the
SAMPEX experimenters.
[email protected]
|
719.28 | September 1-7 | MTWAIN::KLAES | No Guts, No Galaxy | Fri Sep 09 1994 11:24 | 138 |
| From: US1RMC::"[email protected]" "Steve Andrews" 9-SEP-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 9/1/1994
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :Sept. 1, 1994 to Sept. 7, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
-----------------------------------
EUVE continues to perform very well on orbit. A two week
Navigation Certification Test with STGT which began on the 26th,
continues this week.
SPACECRAFT STATUS
EUVE's Explorer Platform performed nominally this past week. The
battery charging strategy was returned to use of 3-amp constant
current mode (CCM) at orbit dawn. This had been discontinued due
to the poor solar angle with the arrays during the observation of
Jupiter, which ended on August 12. In coordination with the GSFC
Space Power Applications Branch, additional data (between the 12th
and Sept. 1) had been collected on battery performance without CCM.
The EUVE Flight Operations Team continues to participate in the
STGT Navigation Certification Test. Problems persist with EUVE's
I-channel data through STGT, and are being troubleshot by the
Space Network.
INSTRUMENT STATUS
No problems or incidents reported.
UPCOMING EVENTS
EUVE begins the week observing the star CF-TUC. Later in the week
we will slew to perform a dithered observation of EUVE-J0030-63.
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
-----------------------------
The Compton observatory is in normal pointing mode. The +Z axis
was oriented toward the Vela Pulsar science target on August 31.
The next attitude maneuver is planned for September 20 to 3C 317,
the last science target in the phase 3 viewing plan.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from August 23 through August 28 was 83 percent, including
19 percent from the GRO Remote Terminal System (GRTS).
Additional coverage of 4 percent was achieved with the BATSE 1
Kbps subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 23 degrees. The
32 Kbps coverage on August 30 was 83 percent, including 20
percent from the GRTS, plus 4 percent of the BATSE 1 Kbps data.
The high-gain antenna boom angle was 17 degrees on that day.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 08/29/94 through 09/03/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 08/29/94 through 09/03/94, 6 real
time passes were supported with 25 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
Observations of Venus took place on Aug 29 and Aug 30.
These were power negative observations at a Sun angle
of 134 degrees using Santiago as a backup station. All
operations were nominal and proceeded without incident.
OPERATIONS
A total of 116 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 61
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 55
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .0 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6720 minutes
Effectiveness. . . . . . . . . . . . . . . . 100%
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
---------------------------------------
The ERBS spacecraft continues to acquire science data from both
the ERBE-NS and the SAGE instruments. The battery and other spacecraft
operations remain stable. ERBS will switch some of its operations to
the new STGT (Second TDRS Ground Terminal) in about one month.
SAMPEX - J. Williamson
-----------------------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
spacecraft is presently in a three week period of continuous
sun-light. The SAMPEX Control Center will reduce its staffing to
only 12 hours a day (7AM - 7PM) from the present 16 hours a day.
This change will require that one of the two recorder dumps a day
be taken in the "blind", i.e. without any Control Center support.
Tests with the Wallops station and PACOR will begin this week.
The purpose of this change is to allow more time for the SWAS and
FAST testing and training which will be operated out of the same
Control Center as SAMPEX.
[email protected]
|
719.29 | September 8-14 | MTWAIN::KLAES | No Guts, No Galaxy | Thu Sep 15 1994 17:00 | 176 |
| From: US1RMC::"[email protected]" "Steve Andrews" 15-SEP-1994
To: [email protected]
CC:
Subj: Orbiting Satellites, 9/8/94 - 9/14/94
CODE :602
INDIVIDUAL:Paul J. Pashby
TITLE :Orbiting Satellites Project
WEEK OF :September 8 to September 14, 1994
EXTREME ULTRAVIOLET EXPLORER (EUVE) - K. Hartnett
-----------------------------------
EUVE apparently suffered the loss of half (TU-1) of its B side
tape recorder on 9/11 which resulted in a small nonrecoverable
data loss. TU-2 shows no signs of stress and continues to be
used. Tape recorder A is fully functional. A problem has also
developed in the data from EUVE payload detector #4. Further
investigation of both problems is underway. EUVE began an
observation of the star FK-2QR on the 10th which will continue
until the 16th. Observations of the Moon and HD220657 are
scheduled to follow.
SPACECRAFT STATUS
There was an apparent hardware failure on tape recorder B, TU-1
on 9/11. Excessive servo errors and missing data characterized
the scheduled 1300Z playback. Subsequent attempts to redump and
reposition the tape recorder during the pass were all
unsuccessful. A nonrecoverable loss of approximately one orbit of
night data from type II target EUVE-J0030 was incurred. EUVE's
two tape recorders (A&B) are comprised of two mechanical
Transport Units (TU) each (TU-1 and TU-2) which share a common
electronics module. TU-2 of tape recorder B has shown no
anomalous behavior and continues to be used. Until the failure
mode of TU-1 is more thoroughly investigated, and as a precautionary
step, TU-2 will be played back only at the 512kb (not 1Mb) rate.
Earlier in the week this same target suffered less than optimal data
coverage at the beginning of its observation because of a coordinate
error in the UCB-provided target science plan. A slew to correct the
minor deviation was executed on 9/8.
In cooperation with the Code 700 GSFC Experimenter, power cycling and
control bus switching commands were also sent to the Environmental
Verification Experiment on the Explorer Platform (EVEEP) on 9/8.
Performance of the EVEEP module has been anomalous since mid July.
The full present status of the experiment is unknown at this time.
A report should be forthcoming this week.
The STGT Navigation Certification test ended on the 9th. EUVE Flight
Controllers have been kept busy documenting the many occurrences of
degraded I-channel (realtime) data following T/R playback
reconfiguration through STGT. Space Network personnel are
investigating the cause. Next week (9/20) a month long test involving
JPL's DSN sites and the Explorer Platform TONS experiment will
commence. No impact to EUVE science is anticipated.
INSTRUMENT STATUS
Spikes were observed on the detector #4 pulse height distribution
on 9/8. Initial analysis indicates that the spikes are at
discreet pulse heights and the same location, implying that the
problem is in the Telemetry Interface (TIF) #4. TIF #4 was reset
and power cycled with no effect. The problem is still under
investigation.
UPCOMING EVENTS
EUVE's targets for the week are FK_AQR, the Moon, and HD220657.
INTERNATIONAL COMETARY EXPLORER (ICE) - G. Bullock
-------------------------------------
Support of the ICE spacecraft continues on a routine basis.
During the week of 09/05/94 through 09/11/94, 2 real time
passes were supported with no commands transmitted.
INTERPLANETARY MONITORING PLATFORM (IMP) - G. Bullock
----------------------------------------
Support of the IMP-8 spacecraft continues on a routine
basis. During the week of 09/05/94 through 09/11/94, 7 real
time passes were supported with 17 commands transmitted.
INTERNATIONAL ULTRAVIOLET EXPLORER (IUE) - G. Bullock
----------------------------------------
HIGHLIGHTS
The ISLU unit for emergency operations with VILSPA,
featuring a new permanent direct data line, was tested
on two occasions this week and has proven to be much
superior to the old system of establishing a modem line
in both setup time involved and data flow speed.
A nine minute loss of science time was realized on
September 8 when the primary ground computer crashed
and a switch to the backup was necessary.
OPERATIONS
A total of 125 images were taken during the week:
Long Wavelength Primary (LWP) . . . . . . . . 62
Long Wavelength Redundant (LWR) . . . . . . . 0
Short Wavelength Primary (SWP) . . . . . . . 63
OPERATIONAL EFFECTIVENESS
Total Time Available : . . . . . . . 6720 minutes
Minus losses due to:
Engineering tests. . . . . . . . . . . .0 minutes
Spacecraft Problems. . . . . . . . . . .0 minutes
Ground System Problems . . . . . . . . .9 minutes
Operations Problems. . . . . . . . . . .0 minutes
Natural Phenomena. . . . . . . . . . . .0 minutes
Time used in Science Data Ops: . . . 6711 minutes
Effectiveness. . . . . . . . . . . . . . . .99.9%
COMPTON GAMMA RAY OBSERVATORY (GRO) - R. Wilson
----------------------------------
The Compton observatory is in normal pointing mode. The +Z axis
has been oriented toward the Vela Pulsar science target since
August 31. The next attitude maneuver is planned for September
20 to 3C 317, the last science target in the phase 3 viewing plan.
Based on the number of packets of data delivered to the
instrument teams from the Packet Processor facility, the
percentage of full-time coverage achieved with 32 Kbps real-time
data from September 1 through September 6 was 82 percent,
including 17 percent from the GRO Remote Terminal System (GRTS).
Additional coverage of 4 percent was achieved with the BATSE 1
Kbps subset of data during this time. The high-gain antenna boom
angle with respect to the equatorial plane was 47 degrees.
EARTH RADIATION BUDGET SATELLITE (ERBS) - J. Williamson
--------------------------------------
The ERBS spacecraft continues to acquire science data from both
the ERBE-NS and the SAGE instruments, and the battery and other
spacecraft operations remain stable. ERBS will switch some of its
operations to the new STGT (Second TDRS Ground Terminal) in about
three weeks.
SAMPEX - J. Williamson
-------
SAMPEX continues to acquire all science and engineering data
without any significant problems or operational errors. The
spacecraft is presently in a three week period of continuous
sun-light. Starting next week, the SAMPEX Control Center will
reduce its staffing to only 12 hours a day (7AM - 7PM) from the
present 16 hours a day. This change will require that one of the
two recorder dumps a day be taken in the "blind", i.e. without
any Control Center support. A number of tests with the Wallops
station and PACOR were successfully performed this past week. The
purpose of this change is to allow more time for the SWAS and
FAST testing and training. The SWAS and FAST missions will be
launched and operated next summer by a combined SAMPEX/SWAS/FAST
Operations Team from the present SMEX Control Center.
[email protected]
|
719.30 | Geminis at GSFC | skylab.zko.dec.com::FISHER | How may I be honest with you today?-Tuvok | Tue Aug 01 1995 15:04 | 29 |
| I got a chance to go to the visitor's center at GSFC last week. To add
to something that .0 did not mention: Of the two Gemini spacecraft,
one is a mockup, the other is Gemini XII. The mockup is mildly
interesting since you can sit in it and get an idea how much space
there is. G12 is really neat, though. It is positioned such that you
can really see it much better than I have seen any other "real"
spacecraft. The right-seat door is open, and you can see all the
controls easily (there is a platform that you get up on to be at the
right level). You are also at a perfect spot to see the heat shield
and the VERY small amount of space between the astronaut's backs and
the shield. Amazing to think of all those thousands of degrees of
plasma flowing by just 6 inches away.
Another interesting thing that you could see easily is the parachute
end. There was a metal ring, apparently a fairly solid piece of stuff,
which I assume the chutes where actually hooked to in some fashion.
Also there were two pyro guillitines which had obviously cut through a
bundle of wires that on the spacecraft end were hooked to a thing
marked something like "pyrotechnic initiator". I assume that these
wires may have run to either the pyros that deployed the chutes, or
possibly the pyros that unfurled them(???did this happen on Gemini?)
and that the bundles are cut on landing as part of releasing the
chutes. Cool stuff to see anyway.
As .0 said, interesting if you are in the area, but not a LOT to see
for a special trip.
Burns
|